### v1

# Mitigate Language Priors in Large Vision-Language Models by Cross-Images Contrastive Decoding

Abstract
Language priors constitute one of the primary causes of hallucinations in Large Vision-Language Models (LVLMs), driving the models to generate linguistically plausible yet visually inconsistent content. The language priors in LVLMs originate from the linguistic knowledge inherited from their pre-trained Large Language Model (LLM) backbone. Consequently, this characteristic is an intrinsic property of the model that remains independent of visual inputs. Inspired by the finding that language priors are consistent across images, we propose Cross-Image Contrastive Decoding (CICD), a simple yet effective training-free method to alleviate language priors in LVLMs. CICD first identifies essential and detrimental priors, and then employs contrastive decoding to eliminate the detrimental ones. This approach simultaneously prevents LVLMs from generating hallucinated content while maintaining textual fluency and coherence. Furthermore, the limited information overlap between images helps prevent visual information loss during contrastive decoding. We validate the effectiveness of CICD on four benchmarks with six LVLMs. Our experiments demonstrate that CICD performs remarkably well in mitigating language priors, especially in the image captioning task, where such priors are most pronounced. Code will be released once accepted.
Index Terms:
Large Vision-Language Model, Multi-Model Hallucination, Language Priors
I Introduction
In recent years, the remarkable success of Large Language Models (LLMs) [1; 2; 3; 4] has provided a viable technological pathway toward Artificial General Intelligence (AGI). Numerous studies [5; 6; 7; 8] have extended these models by incorporating visual inputs, resulting in Large Vision-Language Models (LVLMs) capable of visual comprehension. These models perform various tasks using information from visual inputs, such as image captioning [9; 10; 11; 12], visual question answering [13; 14; 15], and visual grounding [16; 17; 18]. Despite their fluency, LVLMs are prone to hallucinations [19; 20], generating text that contradicts visual evidence.
Hallucinations arise from various factors, most directly from language priors [12; 21; 22; 23], which cause LVLMs to struggle with accurately expressing their understanding of visual information. Since LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge embedded in the LLMs, resulting in responses that reflect both visual cues and language priors. For example, when describing an office scene, an LVLM may incorrectly generate the word “computer” even if no such object is present in the image.
There are many approaches to alleviating language priors. Some works [24; 25; 12] construct specialized training datasets to reduce models’ reliance on language priors. However, the rapid iteration of model backbones limits the practical applicability of such training-based methods. Recently, a series of studies [21; 22; 23; 26] have used the contrastive decoding algorithm [27] to mitigate language priors, achieving remarkable effectiveness. Specifically, these works construct modified contexts by either perturbing [21; 22] or enhancing [23; 26] the visual features from the original context, and then apply contrastive decoding between the original and modified contexts (i.e., subtracting one logit distribution from the other) to reduce language priors.
These methods are sufficiently simple, training-free, and exhibit excellent generalizability. However, since the modified context is derived from the original context, it retains partial visual information, which can be degraded during contrastive decoding. Moreover, the perturbations introduced may corrupt the model’s features, leading to information distortion.
Figure 1: A case of language prior. Adding the same prefix when captioning another image will leads to the same language priors.
The language priors are derived from LLM backbones and are unrelated to the input image. In other words, the language priors remain consistent across different images, as shown in Fig.1. Inspired by this observation, we propose Cross-Images Contrastive Decoding (CICD), a simple yet highly effective method for mitigating language priors in LVLMs. Concretely, we replace the input image with a different image to create a new context that shares the same language priors. We then subtract this logit distribution from the original one to alleviate the influence of language priors, while preserving the original visual information. Since different images contain limited overlapping visual content, our method effectively prevents visual information loss.
Figure 2: Essential priors and detrimental priors. We random select 500 images from the MSCOCO dataset and visualize LLaVA-1.5’s output by t-SNE algorithm [28]. w/ Image represents regular outputs. w/o Image represents the outputs after removing input image. w/ Neg. Image represents replace the input image by the most dissimilar image in the dataset retrieved by CLIP [29].
Figure 3: A case of JS divergence between the logit distribution of two images.
While language priors contribute to hallucinated content in LVLMs, they also play a role in enhancing output fluency. Therefore, we categorize language priors into two types: (1) essential priors, which are necessary for maintaining text fluency, and (2) detrimental priors, which introduce vision-irrelevant information. As shown in Fig. 2, essential priors demonstrate high consistency across different images, whereas detrimental priors exhibit substantial divergence when generating image-related content. Based on this phenomenon, we distinguish between essential and detrimental priors using cross-image consistency, as illustrated in Fig.3. We observe a distinct divergence in cross-image consistency between essential and detrimental priors. We compute the Jensen-Shannon (JS) divergence between the two logit distributions, and set a threshold to identify essential priors that should be retained.
We validate the effectiveness of our method, CICD, on four benchmarks: CHAIR [10], POPE [14], AMBER [30], and DetailCaps [11]. To evaluate the generalizability of CICD, we conduct experiments on six LVLMs: InstructBLIP [6], LLaVA-1.5 [31], LLaVA-Next [32], Qwen-VL-Chat [7], Qwen2-VL [33], and Qwen2.5-VL [8]. Experimental results demonstrate that CICD effectively mitigates language priors in LVLMs, thereby reducing hallucinations. This improvement is particularly notable in the image captioning task, which typically suffers from stronger language priors due to its longer textual context.
In summary, our main contributions are as follows:
  • We conduct a thorough investigation of language priors and demonstrate that they are intrinsic, model-derived properties independent of visual inputs. Furthermore, we show that consistency across different images can be used to distinguish essential priors from detrimental ones.
  • We propose CICD (Cross-Images Contrastive Decoding), which leverages both the consistency of language priors across images and the divergence of inter-image visual features to mitigate linguistic priors while preserving visual information.
  • Extensive experiments across four benchmarks and six model backbones demonstrate the superiority of our method in alleviating language priors, particularly in the image captioning task.
II Related Work
II-A Large Vision-Language Models
Typically, Large Vision-Language Models (LVLMs) [6; 31; 7; 34] consist of a visual encoder [29; 35] and a Large Language Model (LLM) backbone [2; 3]. The visual encoder extracts visual features, which are then aligned with text tokens through a fusion module to form the input context for the LLM. Through multi-stage training, including modality alignment and instruction tuning [5], LVLMs acquire the ability to perform various multimodal tasks such as image captioning [11], visual question answering [15], and visual grounding [18]. Despite significant advancements, LVLMs still suffer from severe hallucination problems, often generating content that contradicts the visual input.
II-B Hallucination in LVLMs
Originally, hallucinations [36; 37] were defined as undesirable outputs or non-factual information. In the multimodal domain, the study of hallucinations primarily focuses on cross-modal consistency [19; 20], requiring that LVLM-generated content remains faithful to the visual inputs. Multiple factors contribute to hallucinations in LVLMs, including biased visual attention [38; 39], attention misallocation [40; 41; 42], and language priors [21; 23; 22]. Among these, language priors are the most direct cause, making it difficult for LVLMs to accurately express their understanding of visual information.
II-C Language Priors
As LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge embedded in LLMs, which are referred to as language priors [24; 25]. Language priors specifically manifest when LVLMs generate content that is linguistically plausible but inconsistent with visual evidence. Some works [12; 43; 44] meticulously design specialized training data to mitigate this issue. However, such training-based approaches exhibit limited generalizability and cannot be easily transferred to broader application scenarios. Recently, a series of training-free approaches have directly reduced language priors using contrastive decoding [27; 45], achieving remarkable performance. These methods construct an alternative logit distribution on top of the original one through techniques such as masking the image [21], perturbing the instruction [22], augmenting the vision input [23], or performing cross-modal conversion [26]. During decoding, the two logit distributions are contrasted to eliminate language priors.
While these methods effectively mitigate linguistic priors, they inevitably incur visual information loss during contrastive decoding due to information overlap, where both distributions share partial visual features. Through a comprehensive analysis of language priors, we identify their vision-independent nature. Based on this insight, we propose the Cross-Images Contrastive Decoding (CICD) method, which eliminates language priors by leveraging their consistency across different images, while preserving visual information through inter-image variation.
III Preliminary and Motivation
LVLMs retain the core architecture of LLMs while augmenting it with a vision encoder for visual feature extraction and contextual integration. The input image is encoded into visual tokens and inserted into the textual context. After forward propagation, LVLMs project the outputs into a logit distribution to compute conditional probabilities and generate responses following an autoregressive paradigm:
$y_{t}\sim p_{\theta}(y_{t}\mid v,x,y_{<t})=\mathrm{softmax}\left[\mathrm{logit}_{\theta}(y_{t}\mid v,x,y{<t})\right]$ (1)
where $v$ and $x$ denote the visual input and textual input, $y_{t}$ denotes the token generated at the $t$-th step, and $y_{<t}=[y_{1},y_{2},\cdots,y_{t-1}]$ is the sequence of tokens generated before $y_{t}$. $\mathrm{logit}_{\theta}$ represents the logit distribution formulated by the LVLM with parameters $\theta$, and is normalized into conditional probabilities $p_{\theta}$.
Since LVLMs are fine-tuned from pre-trained LLMs to enhance fluent generation capabilities, they inherently retain prior knowledge that is independent of visual inputs, i.e., language priors. Concretely, tokens ranked higher in the LVLM’s distribution $\mathrm{logit}_{\theta}$ may be semantically plausible but objectively inconsistent with visual evidence. This phenomenon results in hallucinations, where the generated outputs are linguistically fluent yet visually contradictory.
Language priors can be directly suppressed by downweighting the corresponding tokens in the logit distribution. However, since these tokens cannot be explicitly identified, contrastive decoding [27; 21] has emerged as a widely adopted solution to mitigate language priors. This method constructs an auxiliary distribution with enhanced language priors, and then computes a new distribution by contrasting the original and auxiliary distributions:
$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,x,y_{<t})=$ $(1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$ (2)
$-\alpha\,\mathrm{logit}_{C}$
where $\mathrm{logit}_{C}$ is the auxiliary distribution, and $\mathrm{logit}_{\mathrm{CD}}$ is the final distribution used for next-token prediction.
The core principle of contrastive decoding lies in constructing a logit distribution that preserves consistent language priors while excluding consistent visual information. This ensures that visual information is not inadvertently removed along with language priors during contrastive decoding. However, since LVLMs are fine-tuned using multimodal data, removing the visual input distorts the resulting logit distribution, leading to inconsistent language priors (corresponding experiments are presented in Sec.V-G).
We demonstrate that language priors are inherent characteristics of LVLMs and remain consistent across different images, as shown in Fig.1. An intuitive approach is to use other images to construct a logit distribution that shares the same language priors but contains different visual information. Accordingly, Eq.2 is modified as:
$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,v^{\prime},x,y_{<t})=$ $(1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$ (3)
$-\alpha\,\mathrm{logit}_{\theta}(y_{t}\mid v^{\prime},x,y_{<t})$
where $v\textquoteright$ ’ is a different image from $v$. $\mathrm{logit}_{\theta}(y_{t}\mid v,x,y{<t})$ encodes both the visual information $V$ from image $v$ and the language priors $T$, which we denote as $V\circ T$. The effectiveness of contrastive decoding can be formulated as:
[…]
where $V\textquoteright$ ’ denotes the visual information of image $v\textquoteright$ ’, and $I(V;V\textquoteright)$ ’ ) denotes the consistency between $V$ and $V\textquoteright$ ’. This approach achieves two objectives simultaneously: (1) precise elimination of language priors, and (2) full preservation of visual information integrity.
IV Method
Leveraging the cross-image consistency of language priors, we propose CICD (Cross-Images Contrastive Decoding), which effectively mitigates language priors while preventing visual information loss. In addition, we retain essential linguistic priors to ensure fluent and natural generation. As illustrated in Fig. 4, CICD comprises three key steps: (1) retrieving a distinct image, (2) identifying detrimental language priors, and (3) removing these priors through contrastive decoding.
Figure 4: The input image and another distinct image are fed into LVLM respectively with identical text-prefix. We calculate the JS discrepancy between the two logit distributions to discriminate essential priors and detrimental priors. For detrimental priors, we perform contrastive decoding to alleviate language priors, while applying regular decoding for essential priors. The final output of this case is shown in the right part.
IV-A Image Retrieval
To prevent visual information loss during contrastive decoding, we select an image that has limited information overlap with the input image. Since visual features are extracted by the LVLM’s vision encoder, an effective strategy is to measure the similarity between two images using the vision encoder to estimate their information overlap. In other words, the image with the lowest similarity to the input image is selected as the reference image. For better generalizability, we adopt the pre-trained CLIP [29] as the retriever. We vectorize all images in the dataset in advance, and during inference, retrieve the one with the minimum cosine similarity to the input image.
Although vector retrieval is time-efficient given the computational cost of LVLMs, practical deployment still requires loading an additional retriever and maintaining a vector database. As illustrated in Fig.1, we observe that language priors primarily manifest in content words such as objects, attributes, and relations. These visual elements tend to exhibit minimal information overlap across natural images. As a result, randomly selecting an image can be an effective and lightweight alternative in practical implementations. We further discuss this in Sec.V-G.
To ensure reproducibility, we adopt the retrieval-based approach for deterministic image selection throughout our experiments.
IV-B Identifying Detrimental Priors
Although language priors can lead LVLMs to generate hallucinated content, they are still essential for producing fluent and coherent text. Therefore, we categorize language priors into two types: essential priors and detrimental priors, where only the latter need to be eliminated. We define essential priors as purely linguistic attributes that are entirely independent of visual information. Specifically, for essential priors, the logit distributions produced by LVLMs should remain highly consistent across diverse visual inputs. This intrinsic property enables a clear distinction between essential and detrimental priors.
Figure 5: 500 random images from the MSCOCO dataset captioned by LLaVA-1.5. For each next-token prediction, we statistic the JS divergence (JSD) between two logit distributions produced by different images.
We employ Jensen-Shannon (JS) divergence to measure the consistency between two logit distributions produced by different images:
$\mathcal{D}_{\mathrm{JS}}=\mathrm{JS}\left(p_{\theta}(y_{t}\mid v,x,y_{<t})||p_{\theta}(y_{t}\mid v\textquoteright,x,y_{<t})\right)$ (5)
where $p_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq.1. The statistical results are visualized in Fig.5. We observe that the JSD values can be clearly divided into two distinct ranges: $\log_{10}(\mathcal{D}{\mathrm{JS}})>-4$ and $\log_{10}(\mathcal{D}_{\mathrm{JS}})\leq-4$. Larger JSD values correspond to detrimental priors, while smaller values indicate essential priors. Based on this observation, we set a threshold of $\gamma=-4$ to distinguish between the two types of priors.
IV-C Cross-Images Contrastive Decoding
The model performs regular decoding when essential priors are detected, and applies contrastive decoding to eliminate detrimental priors. For detrimental priors, we hypothesize that smaller distributional differences indicate stronger detrimental effects. Consequently, we define the contrastive coefficient $\alpha$ in Eq. 3 as:
$\alpha=1-\log_{10}(\mathcal{D}_{\mathrm{JS}})$ (6)
Based on the observed density of $\mathcal{D}_{\mathrm{JS}}$ values, we clip $\alpha$ to the interval $[1,3]$ to ensure numerical stability.
The final logit distribution is defined as follows:
$\mathrm{logit}_{CICD}(y_{t}\mid v,v^{\prime},x,y_{<t})=$ (7)
$\begin{cases}\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}\leq\gamma\\ \mathrm{logit}_{CD}(y_{t}\mid v,v^{\prime},x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}>\gamma\end{cases}$
where $\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq. 1, and $\mathrm{logit}_{CD}(y_{t}\mid v,v^{\prime},x,y_{<t})$ is defined in Eq. 3.
V Experiment
V-A Benchmarks
To evaluate the effectiveness of our method in eliminating language priors, we employ four multimodal hallucination benchmarks: two generative benchmarks (CHAIR [10] and DetailCaps [11]), one discriminative benchmark (POPE [14]), and one hybrid benchmark (AMBER [30]).
CHAIR assesses the proportion of hallucinated objects, which are generated by the model but not present in the reference annotations. Following prior works, we randomly select 500 images from the MSCOCO [46] dataset as the test set. We additionally select another 500 images to construct a validation set for hyperparameter tuning.
DetailCaps is a fine-grained image captioning benchmark, accompanied by ground-truth detail captions generated by GPT-4V [47], Gemini-1.5-Pro [48], and GPT-4o [49] for evaluation. It contains 4,870 images from multiple datasets, and we use a subset of 700 images from MSCOCO to conduct our experiments.
POPE is a widely adopted benchmark for evaluating object hallucinations by prompting LVLMs to identify whether a specific object is present in the image. It consists of 1,500 images, including 1,000 from MSCOCO and 500 from GQA [50].
AMBER combines generative and discriminative tasks, and is evaluated on a curated set of 1,004 images. In addition to image captioning, it includes 14,216 questions designed to assess hallucinations in object, attribute, and relation recognition.
V-B Evaluated LVLMs
We evaluate our method on six widely-used LVLMs: InstructBLIP [6], LLaVA-1.5 [31], LLaVA-Next [32], Qwen-VL-Chat [7], Qwen2-VL [33], and Qwen2.5-VL [8]. InstructBLIP and LLaVA-1.5 are based on Vicuna-7B [3] as their language backbone, while LLaVA-Next is built upon LLaMA3. The Qwen-VL series models, including Qwen2-VL and Qwen2.5-VL, are developed based on the QwenLM framework. Regarding vision encoders, InstructBLIP, LLaVA-1.5, LLaVA-Next, and Qwen-VL-Chat all use the pretrained CLIP model [29], whereas Qwen2-VL and Qwen2.5-VL employ a fine-tuned ViT [35]. All models are implemented at the 8B parameter scale.
TABLE I: Results on POPE. Acc. and F1 represent the average accuracy and F1-score across three datasets (MSCOCO, A-OKVQA, and GQA), respectively. The bolded results denote the best in each setting and the underlined results denote the second-best.
Setting Method LLaVA-1.5 InstructBLIP Qwen-VL-Chat
Acc. $\uparrow$ F1$\uparrow$ Acc. $\uparrow$ F1$\uparrow$ Acc. $\uparrow$ F1$\uparrow$
Random Regular 83.47 82.25 80.49 81.10 82.32 79.35
VCD 85.30 84.59 81.75 82.06 84.74 82.80
ICD 84.81 83.44 80.95 81.59 83.10 80.52
IBD 85.36 84.28 74.82 77.87 83.07 80.23
DeFG 86.68 85.71 84.35 84.50 84.04 81.49
CICD 87.85 87.77 84.00 84.89 87.54 86.37
Popular Regular 79.56 78.99 76.59 78.16 81.11 78.34
VCD 80.81 80.84 77.22 78.60 82.95 81.17
ICD 81.44 80.53 76.82 78.48 81.26 78.87
IBD 81.68 81.13 69.35 74.31 82.20 79.59
DeFG 82.96 82.46 79.25 80.46 83.02 80.57
CICD 84.83 85.18 81.01 82.53 85.94 84.90
Adversarial Regular 76.95 77.13 72.89 75.58 78.23 75.72
VCD 77.27 78.19 74.03 76.39 80.37 78.96
ICD 78.07 77.98 73.19 75.91 79.03 77.05
IBD 78.26 78.52 66.60 72.63 79.68 77.36
DeFG 79.40 79.69 75.19 77.52 80.42 78.24
CICD 78.60 80.30 75.55 78.63 81.94 81.37
V-C Baselines
We compare our method with several training-free baselines that specifically target language priors. VCD [21] introduces noise into the input image to suppress language priors, and then applies contrastive decoding for their elimination. ICD [22] constructs negative logits using adversarial instructions and enhances positive signals through contrastive decoding. IBD [23] strengthens the model’s attention to visual features, thereby increasing the influence of image information during decoding. DeFG [26] converts the generated caption into an image and re-generates the caption from that image, leveraging the preserved visual content to reduce hallucinations.
V-D Implementation Details
We use sampling-based decoding for next-token prediction with default settings. Following [21], we incorporate adaptive plausibility constraints into the decoding process. We use the MSCOCO 2014 validation set as the image retrieval corpus, which contains 40,504 images. For retrieval, we adopt CLIP111openai/clip-vit-large-patch14-336 as the feature extractor and perform similarity-based retrieval using precomputed image embeddings. All experiments are conducted on a single 24GB NVIDIA RTX 3090 GPU.
V-E Main Results
Results on POPE
POPE is a discriminative benchmark that incorporates three distinct datasets: MSCOCO, A-OKVQA, and GQA. Each dataset uses three different negative sampling settings: Random, Popular, and Adversarial. We conducted a comprehensive comparison between our method and other baselines across three distinct LVLMs, as detailed in Tab. I. For streamlined presentation, we report the average performance across all three datasets as the final result.
Our method demonstrates superior overall performance, particularly in the most critical metric, achieving the highest F1 scores across all experimental settings. This suggests that we can utilize cross-image consistency to effectively alleviate language priors and promote LVLMs to convey faithful visual information. Additionally, our method demonstrates consistently superior performance across diverse LVLMs, evidencing remarkable generalization capability. As a training-free approach, it offers immediate deployability in practical applications. Another key finding reveals that our method achieves significantly enhanced performance in the Popular setting, demonstrating an average F1 score improvement of 2.84 over the second-best approach. The Popular setting uses the most frequent objects to construct negative samples, which makes LVLMs more susceptible to language priors. The most substantial performance improvements achieved by our method under this setting further demonstrate its exceptional efficacy in mitigating language priors.
TABLE II: Results on the test set of CHAIR. MaxLen=512 indicates setting the maximum generation length of 512 tokens for the model. The bolded results denote the best in each setting and the underlined results denote the second-best.
Setting Method LLaVA-1.5 InstructBLIP
[…]
MaxLen=512 Regular 54.6 16.4 72.6 106.8 62.6 19.5 66.9 113.6
VCD 59.8 17.8 75.6 104 64.8 18.8 71.9 108.6
ICD 57.0 15.0 74.6 103.3 59.0 17.1 69.2 108.5
IBD 57.6 16.5 74.2 104.3 57.6 15.7 70.8 118.5
DeFG 57.4 16.3 75.5 102.8 59.0 17.7 71.5 108.9
CICD 43.8 11.7 75.0 103.3 49.8 13.7 70.3 98.8
MaxLen=64 Regular 24.4 8.9 56.6 53.7 35.6 13.2 56.4 54.8
VCD 25.0 8.3 59 54.0 32.2 10.3 60.6 55.2
ICD 23.2 8.1 58.4 54.4 29.8 9.8 60.6 55.7
IBD 21.2 6.9 58.8 54.5 27.8 9.2 60.3 56.1
DeFG 22.4 7.2 58.2 54.4 32.4 11 59.4 56.0
CICD 18.0 6.1 59.6 53.3 23.8 7.7 62.2 54.8
TABLE III: Results on AMBER. We set the maximun generation length to 512 for the models in this benchmark. AMBER metirc is calculated by $(100-CHAIR+F1)/2$. The bolded results denote the best in each setting and the underlined results denote the second-best.
Model Method Generative Discriminative
[…]
LLaVA-1.5 Regular 11.6 49.7 47.7 4.4 67.4 83.9 61.8 71.2 79.80
VCD 9.8 51.2 43.8 4.4 68.1 85.1 61.0 71.1 80.65
ICD 8.8 51.2 38.7 4.1 70.3 86.0 64.1 73.4 82.30
IBD 9.8 50.5 42.2 4.4 69.2 86.3 62.1 72.2 81.20
DeFG 9.1 50.7 39.9 4.1 70.2 86.8 63.0 73.0 81.95
CICD 6.6 52.7 34.8 2.2 71.1 89.4 61.9 73.1 83.25
InstructBLIP Regular 12.4 51.9 52.4 5.0 68.3 78.6 71.7 75 81.30
VCD 9.9 54.0 44.6 4.2 70.6 80.6 72.7 76.4 83.25
ICD 9.8 53.9 46.7 5.1 69.9 80.3 70.9 75.3 82.75
IBD 9.0 56.1 45.1 4.6 54.2 86.3 34.2 49.00 70.00
DeFG 9.7 54.1 44.5 5.2 72.1 80.7 75.1 77.8 84.05
CICD 7.1 53.6 35.0 2.3 72.8 85.4 70.1 77.0 84.95
Results on CHAIR
CHAIR is specifically designed to assess object hallucination in the image captioning task, which is one of the most language-prior-susceptible tasks in multimodal research. We evaluate our method in two settings with different maximum generation lengths; the results are shown in Tab. II.
Our method demonstrates superior performance across all experimental conditions, with particularly pronounced improvements observed at longer generation lengths. Longer context lengths amplify the model’s susceptibility to language priors, making these experimental results particularly compelling evidence of our method’s efficacy in mitigating such priors. Compared to baselines, our method’s principal advantage lies in its complete preservation of visual information, which is directly reflected in the superior experimental results. Additionally, we preserve the essential language priors, which are also important in generative tasks.
Results on AMBER
AMBER comprises generative tasks (image captioning) and discriminative tasks (image understanding). The results are shown in Tab. III. We can observe that our method achieves the lowest hallucination frequency in generative tasks while simultaneously attaining state-of-the-art performance on the comprehensive AMBER metric. This superior performance stems from our focused approach to language prior mitigation. This problem is particularly critical for generative tasks, where language priors manifest more severely than in discriminative tasks, explaining our method’s pronounced advantages in image captioning.
TABLE IV: Results on DetailCaps. The bolded results denote the best in each setting and the underlined results denote the second-best.
[…]
LLaVA-1.5 InstructBLIP Qwen-VL-Chat
Regular 52.60 52.99 30.73
VCD 52.91 53.20 31.28
ICD 52.82 53.24 29.70
IBD 52.48 54.14 29.31
DeFG 52.72 53.06 30.21
CICD 55.80 54.20 31.34
Results on DetailCaps
We further validate our method’s effectiveness on the image captioning task, which is more susceptible to language priors. Tab. IV presents the evaluation results on DetailCaps. The CAPTURE metric quantitatively evaluates caption quality through both lexical matching and semantic alignment, measuring the consistency of key content (entities, attributes, and relations) between generated captions and multiple ground-truth references. Our method achieves the best performance across all three evaluated LVLMs, demonstrating its robust capability to mitigate language priors and significantly enhance image captioning performance.
TABLE V: Results on more LVLMs. F1 is the average F1-score across three datasets (MSCOCO, A-OKVQA, and GQA). The bolded results denote the best in each setting.
Model Method CHAIR(MaxLen=512) POPE (F1$\uparrow$)
[…]
LLaVA-next Regular 39.0 10.8 62.1 85.47 83.14 79.87
CICD 37.6 10.4 62.7 88.22 86.29 82.76
Qwen2-VL Regular 53.8 10.0 69.4 88.33 85.82 82.43
CICD 35.0 9.2 60.5 90.67 88.07 84.53
Qwen2.5-VL Regular 41.2 11.5 65.7 87.83 85.66 83.12
CICD 37.2 9.3 60.0 89.95 88.00 84.75
V-F Effectiveness on More LVLMs
To comprehensively demonstrate the effectiveness and generalizability of our method, we select three of the most recent open-source LVLMs to conduct experiments. The results are presented in Tab. V. Our method demonstrates significant performance improvements in all experiments, consistently reducing hallucinations across both generative and discriminative tasks through effective mitigation of language priors. The improvements observed across different LVLMs confirm our method’s practical advantages: an elegantly simple framework coupled with state-of-the-art effectiveness.
TABLE VI: This experiment conducted by LLaVA-1.5. Random Image and Retrieved Image represent the contrastive images selected through random sampling and image retrieval, respectively. w/o Image represents the contrastive distribution is constrcuted by removing the input image.
[…]
CHAIR(MaxLen=512)
Regular 54.6 16.4 72.6
w/o Image 54.6 15.5 76.6
Random Image 46.6 13.2 76.7
Retrieved Image 43.8 11.7 75
CHAIR(MaxLen=64)
Regular 24.4 8.9 56.6
w/o Image 25.4 7.6 61.9
Random Image 21.6 6.9 58.8
Retrieved Image 18.0 6.1 59.6
V-G Divergence Between Images
Contrastive decoding is a method that suppresses consistent features between two distributions while amplifying distinctive ones. To prevent the loss of visual information during contrastive decoding, we select an image with sufficient discrepancy to construct a contrasting distribution. Thus, while maintaining language prior consistency, the divergent visual information helps avoid erroneous debiasing caused by feature overlap. We use CLIP to retrieve images based on cosine similarity, and select the image with the lowest similarity as the contrastive image. We conduct experiments to explore the correlation between similarity and performance, as shown in Fig. 6.
We observe that the similarity between images exhibits no strong correlation with performance. Compared to regular decoding methods, images with varying similarity levels all demonstrate significant improvements. As discussed in Sec.IV-A, the influence of language priors is primarily reflected in the LVLMs’ comprehension of objects, attributes, and relationships. This information exhibits limited overlap across real-world images. We hypothesize that random image selection could achieve comparable effectiveness, with experimental results presented in Tab.VI. Experimental results demonstrate that random image selection can also significantly mitigate language priors in LVLMs. In practical applications, the random selection approach offers greater convenience, substantially enhancing the usability of our method.
In Tab. VI, we also attempted to construct language priors by removing visual inputs, but found this approach less effective compared to eliminating linguistic priors through cross-image consistency. We attribute this limitation to LVLMs being trained on large-scale multimodal data, where removing visual inputs disrupts the model’s intrinsic representations and leads to information distortion. Consequently, language priors in LVLMs cannot be directly constructed, which further substantiates the necessity of leveraging cross-image consistency in our method.
Figure 6: The correlation between similarity and performance. We conduct this experiment on the validation set of CHAIR. Sim-CLIP represents normalized image similarity, measured as the cosine distance between the features of two images extracted by CLIP.
Figure 7: The results on the validation set of CHAIR using LLaVA-1.5. The X-axis represents the value of $\gamma$.
V-H Threshold of Detrimental Priors
While language priors contribute to hallucinations in LVLMs, they also play a critical role in maintaining textual coherence and fluency. We employ JS divergence to distinguish between essential and detrimental priors, where essential priors exhibit high cross-image consistency, reflected in lower JS divergence scores. We establish a threshold of $\gamma=-4$ in Eq.7, based on the statistical distribution of JS divergence scores shown in Fig.5.
A comprehensive analysis is shown in Fig. 7. The experimental results align with the statistical analysis, demonstrating optimal performance at $\gamma=-4$. Another notable finding is that preserving essential priors proves meaningful, as it ensures fluent and coherent generation contexts.
VI Conclusion
In this paper, we focus on the language priors issue in LVLMs. Through a comprehensive investigation, we reveal that language priors are intrinsic characteristics of the model and remain consistent across different images. Based on this observation, we propose Cross-Images Contrastive Decoding (CICD)—a training-free method that effectively mitigates language priors while preserving visual information. Extensive experiments across multiple benchmarks and LVLM architectures validate the effectiveness of CICD, particularly in image captioning tasks where language priors are most pronounced.
The simplicity and efficiency of CICD give it substantial practical potential, enabling synergistic integration with other hallucination mitigation methods to further enhance LVLM performance. This constitutes our primary direction for future work.

### v2

@@
 Abstract
-Language priors constitute one of the primary causes of hallucinations in Large Vision-Language Models (LVLMs), driving the models to generate linguistically plausible yet visually inconsistent content.
-The language priors in LVLMs originate from the linguistic knowledge inherited from their pre-trained Large Language Model (LLM) backbone.
-Consequently, this characteristic is an intrinsic property of the model that remains independent of visual inputs.
-Inspired by the finding that language priors are consistent across images, we propose Cross-Image Contrastive Decoding (CICD), a simple yet effective training-free method to alleviate language priors in LVLMs.
-CICD first identifies essential and detrimental priors, and then employs contrastive decoding to eliminate the detrimental ones.
-This approach simultaneously prevents LVLMs from generating hallucinated content while maintaining textual fluency and coherence.
-Furthermore, the limited information overlap between images helps prevent visual information loss during contrastive decoding.
-We validate the effectiveness of CICD on four benchmarks with six LVLMs.
-Our experiments demonstrate that CICD performs remarkably well in mitigating language priors, especially in the image captioning task, where such priors are most pronounced.
-Code will be released once accepted.
-Index Terms:
-Large Vision-Language Model, Multi-Model Hallucination, Language Priors
-I Introduction
-In recent years, the remarkable success of Large Language Models (LLMs) [1; 2; 3; 4] has provided a viable technological pathway toward Artificial General Intelligence (AGI).
-Numerous studies [5; 6; 7; 8] have extended these models by incorporating visual inputs, resulting in Large Vision-Language Models (LVLMs) capable of visual comprehension.
-These models perform various tasks using information from visual inputs, such as image captioning [9; 10; 11; 12], visual question answering [13; 14; 15], and visual grounding [16; 17; 18].
-Despite their fluency, LVLMs are prone to hallucinations [19; 20], generating text that contradicts visual evidence.
-Hallucinations arise from various factors, most directly from language priors [12; 21; 22; 23], which cause LVLMs to struggle with accurately expressing their understanding of visual information.
-Since LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge embedded in the LLMs, resulting in responses that reflect both visual cues and language priors.
-For example, when describing an office scene, an LVLM may incorrectly generate the word “computer” even if no such object is present in the image.
-There are many approaches to alleviating language priors.
-Some works [24; 25; 12] construct specialized training datasets to reduce models’ reliance on language priors.
-However, the rapid iteration of model backbones limits the practical applicability of such training-based methods.
-Recently, a series of studies [21; 22; 23; 26] have used the contrastive decoding algorithm [27] to mitigate language priors, achieving remarkable effectiveness.
-Specifically, these works construct modified contexts by either perturbing [21; 22] or enhancing [23; 26] the visual features from the original context, and then apply contrastive decoding between the original and modified contexts (i.e., subtracting one logit distribution from the other) to reduce language priors.
-These methods are sufficiently simple, training-free, and exhibit excellent generalizability.
-However, since the modified context is derived from the original context, it retains partial visual information, which can be degraded during contrastive decoding.
-Moreover, the perturbations introduced may corrupt the model’s features, leading to information distortion.
-Figure 1: A case of language prior. Adding the same prefix when captioning another image will leads to the same language priors.
-The language priors are derived from LLM backbones and are unrelated to the input image.
-In other words, the language priors remain consistent across different images, as shown in Fig.1.
-Inspired by this observation, we propose Cross-Images Contrastive Decoding (CICD), a simple yet highly effective method for mitigating language priors in LVLMs.
-Concretely, we replace the input image with a different image to create a new context that shares the same language priors.
-We then subtract this logit distribution from the original one to alleviate the influence of language priors, while preserving the original visual information.
-Since different images contain limited overlapping visual content, our method effectively prevents visual information loss.
-Figure 2: Essential priors and detrimental priors.
-We random select 500 images from the MSCOCO dataset and visualize LLaVA-1.5’s output by t-SNE algorithm [28].
+Language priors are a major cause of hallucinations in Large Vision-Language Models (LVLMs), often leading to text that is linguistically plausible but visually inconsistent.
+Recent work explores contrastive decoding as a training-free solution, but these methods typically construct negative visual contexts from the original image, resulting in visual information loss and distorted distribution.
+Motivated by the observation that language priors stem from the LLM backbone and remain consistent across images, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective training-free method that uses different images to construct negative visual contexts.
+We further analyze the cross-image behavior of language priors and introduce a distinction between essential priors (supporting fluency) and detrimental priors (causing hallucinations), enabling selective suppression.
+By selectively preserving essential priors and suppressing detrimental ones, our method reduces hallucinations while maintaining coherent and fluent language generation.
+Experiments on four benchmarks and six LVLMs across three model families confirm the effectiveness and generalizability of CICD, especially in image captioning, where language priors are particularly pronounced.
+Code will be released upon acceptance.
+Mitigate Language Priors in Large Vision-Language Models by Cross-Images Contrastive Decoding
+Jianfei Zhao1,2, Feng Zhang1, Xin Sun1, Chong Feng1,3,thanks: Corresponding Autor: chongfeng@bit.edu.cn , 1School of Computer Science and Technology, Beijing Institute of Technology, Beijing, China, 2Zhongguancun Academy, Beijing, China, 3Southeast Academy of Information Technology, Beijing Institute of Technology, Fujian, China,
+1 Introduction
+Figure 1: A case of language prior. Adding the same prefix when captioning another image will leads to the same language priors. The model used is LLaVA-1.5.
+Recent advances in Large Language Models (LLMs) Brown et al.
+(2020); Touvron et al.
+(2023); Chiang et al.
+(2023); Bai et al.
+(2023a) have opened promising pathways toward Artificial General Intelligence (AGI).
+Building on these developments, numerous studies Liu et al.
+(2023); Dai et al.
+(2023); Bai et al.
+(2023b, 2025) have extended LLMs by integrating visual inputs, giving rise to Large Vision-Language Models (LVLMs) capable of understanding visual content.
+They are capable of addressing a broad spectrum of vision-language tasks, such as image captioning Jing et al.
+(2024); Rohrbach et al.
+(2018); Ye et al.
+(2025); Chen et al.
+(2025), visual question answering (VQA) Lovenia et al.
+(2024); Li et al.
+(2023b); Yu et al.
+(2024), and visual grounding Rasheed et al.
+(2024); You et al.
+(2024); Lu et al.
+(2024).
+Despite impressive capabilities, LVLMs remain susceptible to hallucinations Li et al.
+(2023c); Wang et al.
+(2023), often generating text that contradicts the visual input.
+These hallucinations stem from various sources, with language priors Chen et al.
+(2025); Leng et al.
+(2024); Wang et al.
+(2024c); Zhu et al.
+(2024b) being the most fundamental underlying cause.
+Since LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge encoded in the LLMs.
+As a result, their outputs are influenced not only by visual evidence but also by pre-learned language patterns.
+This often leads to a tendency to rely on memorized linguistic associations when responding to user queries.
+For instance, when describing an office scene, an LVLM may erroneously mention a "computer" even if no such object is present in the image.
+Figure 2: Essential priors (left) and detrimental priors (right).
+We randomly select 500 images from the MSCOCO dataset and visualize LLaVA-1.5’s output by the t-SNE algorithm van der Maaten and Hinton (2008).
 w/ Image represents regular outputs.
-w/o Image represents the outputs after removing input image.
+w/o Image represents the outputs after removing the input image.
 w/ Neg.
-Image represents replace the input image by the most dissimilar image in the dataset retrieved by CLIP [29].
-Figure 3: A case of JS divergence between the logit distribution of two images.
-While language priors contribute to hallucinated content in LVLMs, they also play a role in enhancing output fluency.
-Therefore, we categorize language priors into two types: (1) essential priors, which are necessary for maintaining text fluency, and (2) detrimental priors, which introduce vision-irrelevant information.
-As shown in Fig.
-2, essential priors demonstrate high consistency across different images, whereas detrimental priors exhibit substantial divergence when generating image-related content.
-Based on this phenomenon, we distinguish between essential and detrimental priors using cross-image consistency, as illustrated in Fig.3.
-We observe a distinct divergence in cross-image consistency between essential and detrimental priors.
-We compute the Jensen-Shannon (JS) divergence between the two logit distributions, and set a threshold to identify essential priors that should be retained.
-We validate the effectiveness of our method, CICD, on four benchmarks: CHAIR [10], POPE [14], AMBER [30], and DetailCaps [11].
-To evaluate the generalizability of CICD, we conduct experiments on six LVLMs: InstructBLIP [6], LLaVA-1.5 [31], LLaVA-Next [32], Qwen-VL-Chat [7], Qwen2-VL [33], and Qwen2.5-VL [8].
-Experimental results demonstrate that CICD effectively mitigates language priors in LVLMs, thereby reducing hallucinations.
-This improvement is particularly notable in the image captioning task, which typically suffers from stronger language priors due to its longer textual context.
+Image represents replacing the input image by the most dissimilar image in the dataset retrieved by CLIP Radford et al.
+(2021).
+Methods for mitigating language priors in LVLMs can generally be classified into two categories.
+The first involves constructing specialized training datasets to reduce the model’s dependence on linguistic biases Zhibo et al.
+(2023); Wu et al.
+(2022); Chen et al.
+(2025), though such approaches often incur substantial training costs.
+In this work, we focus on plug-and-play, training-free methods.
+Many recent studies adopt contrastive decoding Li et al.
+(2023a), which constructs negative contexts to expose and suppress language priors.
+Some approaches directly perturb the input image Leng et al.
+(2024); Wang et al.
+(2024c) to alter the visual context, while others enhance the visual features extracted from the original image Zhu et al.
+(2024b); Zhang et al.
+(2025).
+These methods typically compute the difference between the logit distributions obtained from the original and modified (i.e., negative) contexts to reduce the model’s reliance on language priors.
+Following the paradigm introduced in VCD Leng et al.
+(2024), these methods routinely modify the original image to generate negative visual contexts.
+However, this prevailing design choice raises a critical question: Is constructing negative contexts through perturbations of the original image truly the most effective strategy for inducing and mitigating language priors?
+We argue that the answer is unequivocally negative.
+Constructing negative contexts based on the original image can lead to two main issues: (1) Visual information loss, where the negative visual context retains partial visual information, which may be degraded during contrastive decoding.
+(2) Distorted distribution, where perturbing the model’s input to induce priors may disrupt its internal representations, ultimately resulting in a distorted logit distribution.
+This analysis raises an important question: How can the above issues be avoided when constructing negative contexts?
+In fact, simply taking another image to construct negative context is sufficient.
+This insight is based on the thought that the language priors guiding step-wise token generation in LVLMs originate from the LLM backbone and are largely independent of the input image.
+As shown in Fig.1, LVLM produces highly consistent outputs when given completely different images but the same prefix, which strongly supports our hypothesis.
+Building on this, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective method for mitigating language priors in LVLMs.
+Instead of modifying the original image, CICD constructs the negative context using a different image that shares the same language priors.
+Our method avoids the two key issues present in previous approaches.
+Since the overlap between the original and the new image is limited, our approach avoids introducing critical visual information into the negative context.
+Moreover, it enables the model to reveal intrinsic priors, maintaining the model’s natural inference process without injecting external bias.
+The formal analysis of our method is presented in Sec.
+3.2.
+Figure 3: A case of JS divergence between the logit distributions of two images.
+Focusing on language priors, we further explore whether all such priors should be strictly eliminated.
+We argue that the answer is negative.
+Language priors can be divided into two categories: (1) essential priors, which are necessary for maintaining textual fluency, and (2) detrimental priors, which introduce information unrelated to the visual input.
+Clearly, only the detrimental priors should be removed.
+This raises a critical question: How can we accurately distinguish between the two and selectively eliminate only the detrimental ones?
+As shown in Fig.2, essential priors exhibit high consistency across different images, whereas detrimental priors show substantial divergence when generating image-related content.
+Based on this observation, we propose to distinguish essential from detrimental priors using cross-image consistency, as illustrated in Fig.3.
+Specifically, we compute the Jensen–Shannon (JS) divergence between the logit distributions of different visual contexts, and apply a stable threshold to identify essential priors that should be retained.
+We validate the effectiveness of CICD on four widely used benchmarks: CHAIR Rohrbach et al.
+(2018), POPE Li et al.
+(2023b), AMBER Wang et al.
+(2024a), and DetailCaps Ye et al.
+(2025).
+To assess the generalizability of our approach, we conduct experiments across six LVLMs from three different model families: InstructBLIP Dai et al.
+(2023), LLaVA-1.5 Liu et al.
+(2024c), LLaVA-Next Liu et al.
+(2024d), Qwen-VL-Chat Bai et al.
+(2023b), Qwen2-VL-Instruct Wang et al.
+(2024b), and Qwen2.5-VL-Instruct Bai et al.
+(2025).
+Experimental results show that CICD consistently reduces hallucinations by effectively mitigating language priors in LVLMs, outperforming state-of-the-art methods across all benchmarks and models.
+The improvement is particularly pronounced in the image captioning task, which tends to suffer more from language priors due to its longer textual outputs.
 In summary, our main contributions are as follows:
-  • We conduct a thorough investigation of language priors and demonstrate that they are intrinsic, model-derived properties independent of visual inputs.
-Furthermore, we show that consistency across different images can be used to distinguish essential priors from detrimental ones.
-  • We propose CICD (Cross-Images Contrastive Decoding), which leverages both the consistency of language priors across images and the divergence of inter-image visual features to mitigate linguistic priors while preserving visual information.
-  • Extensive experiments across four benchmarks and six model backbones demonstrate the superiority of our method in alleviating language priors, particularly in the image captioning task.
-II Related Work
-II-A Large Vision-Language Models
-Typically, Large Vision-Language Models (LVLMs) [6; 31; 7; 34] consist of a visual encoder [29; 35] and a Large Language Model (LLM) backbone [2; 3].
-The visual encoder extracts visual features, which are then aligned with text tokens through a fusion module to form the input context for the LLM.
-Through multi-stage training, including modality alignment and instruction tuning [5], LVLMs acquire the ability to perform various multimodal tasks such as image captioning [11], visual question answering [15], and visual grounding [18].
-Despite significant advancements, LVLMs still suffer from severe hallucination problems, often generating content that contradicts the visual input.
-II-B Hallucination in LVLMs
-Originally, hallucinations [36; 37] were defined as undesirable outputs or non-factual information.
-In the multimodal domain, the study of hallucinations primarily focuses on cross-modal consistency [19; 20], requiring that LVLM-generated content remains faithful to the visual inputs.
-Multiple factors contribute to hallucinations in LVLMs, including biased visual attention [38; 39], attention misallocation [40; 41; 42], and language priors [21; 23; 22].
-Among these, language priors are the most direct cause, making it difficult for LVLMs to accurately express their understanding of visual information.
-II-C Language Priors
-As LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge embedded in LLMs, which are referred to as language priors [24; 25].
-Language priors specifically manifest when LVLMs generate content that is linguistically plausible but inconsistent with visual evidence.
-Some works [12; 43; 44] meticulously design specialized training data to mitigate this issue.
-However, such training-based approaches exhibit limited generalizability and cannot be easily transferred to broader application scenarios.
-Recently, a series of training-free approaches have directly reduced language priors using contrastive decoding [27; 45], achieving remarkable performance.
-These methods construct an alternative logit distribution on top of the original one through techniques such as masking the image [21], perturbing the instruction [22], augmenting the vision input [23], or performing cross-modal conversion [26].
-During decoding, the two logit distributions are contrasted to eliminate language priors.
-While these methods effectively mitigate linguistic priors, they inevitably incur visual information loss during contrastive decoding due to information overlap, where both distributions share partial visual features.
-Through a comprehensive analysis of language priors, we identify their vision-independent nature.
-Based on this insight, we propose the Cross-Images Contrastive Decoding (CICD) method, which eliminates language priors by leveraging their consistency across different images, while preserving visual information through inter-image variation.
-III Preliminary and Motivation
-LVLMs retain the core architecture of LLMs while augmenting it with a vision encoder for visual feature extraction and contextual integration.
-The input image is encoded into visual tokens and inserted into the textual context.
-After forward propagation, LVLMs project the outputs into a logit distribution to compute conditional probabilities and generate responses following an autoregressive paradigm:
-$y_{t}\sim p_{\theta}(y_{t}\mid v,x,y_{<t})=\mathrm{softmax}\left[\mathrm{logit}_{\theta}(y_{t}\mid v,x,y{<t})\right]$ (1)
-where $v$ and $x$ denote the visual input and textual input, $y_{t}$ denotes the token generated at the $t$-th step, and $y_{<t}=[y_{1},y_{2},\cdots,y_{t-1}]$ is the sequence of tokens generated before $y_{t}$.
-$\mathrm{logit}_{\theta}$ represents the logit distribution formulated by the LVLM with parameters $\theta$, and is normalized into conditional probabilities $p_{\theta}$.
-Since LVLMs are fine-tuned from pre-trained LLMs to enhance fluent generation capabilities, they inherently retain prior knowledge that is independent of visual inputs, i.e., language priors.
-Concretely, tokens ranked higher in the LVLM’s distribution $\mathrm{logit}_{\theta}$ may be semantically plausible but objectively inconsistent with visual evidence.
-This phenomenon results in hallucinations, where the generated outputs are linguistically fluent yet visually contradictory.
-Language priors can be directly suppressed by downweighting the corresponding tokens in the logit distribution.
-However, since these tokens cannot be explicitly identified, contrastive decoding [27; 21] has emerged as a widely adopted solution to mitigate language priors.
-This method constructs an auxiliary distribution with enhanced language priors, and then computes a new distribution by contrasting the original and auxiliary distributions:
-$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,x,y_{<t})=$ $(1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$ (2)
-$-\alpha\,\mathrm{logit}_{C}$
-where $\mathrm{logit}_{C}$ is the auxiliary distribution, and $\mathrm{logit}_{\mathrm{CD}}$ is the final distribution used for next-token prediction.
-The core principle of contrastive decoding lies in constructing a logit distribution that preserves consistent language priors while excluding consistent visual information.
-This ensures that visual information is not inadvertently removed along with language priors during contrastive decoding.
-However, since LVLMs are fine-tuned using multimodal data, removing the visual input distorts the resulting logit distribution, leading to inconsistent language priors (corresponding experiments are presented in Sec.V-G).
-We demonstrate that language priors are inherent characteristics of LVLMs and remain consistent across different images, as shown in Fig.1.
-An intuitive approach is to use other images to construct a logit distribution that shares the same language priors but contains different visual information.
-Accordingly, Eq.2 is modified as:
-$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,v^{\prime},x,y_{<t})=$ $(1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$ (3)
-$-\alpha\,\mathrm{logit}_{\theta}(y_{t}\mid v^{\prime},x,y_{<t})$
-where $v\textquoteright$ ’ is a different image from $v$.
-$\mathrm{logit}_{\theta}(y_{t}\mid v,x,y{<t})$ encodes both the visual information $V$ from image $v$ and the language priors $T$, which we denote as $V\circ T$.
-The effectiveness of contrastive decoding can be formulated as:
-[…]
-where $V\textquoteright$ ’ denotes the visual information of image $v\textquoteright$ ’, and $I(V;V\textquoteright)$ ’ ) denotes the consistency between $V$ and $V\textquoteright$ ’.
-This approach achieves two objectives simultaneously: (1) precise elimination of language priors, and (2) full preservation of visual information integrity.
-IV Method
+  • We conduct an in-depth analysis of existing contrastive decoding methods and identify their inherent limitations rooted in the reliance on the original image.
+We further verify the consistency of language priors across images and propose a stable approach for distinguishing essential priors from detrimental ones.
+  • Based on this analysis, we propose CICD (Cross-Images Contrastive Decoding), which leverages retrieved images to induce language priors while selectively preserving essential priors during the mitigation process.
+  • Extensive experiments on multiple benchmarks and LVLMs demonstrate the superiority of our method in alleviating language priors, particularly in the image captioning task.
+2 Related Work
+Due to space limitations, we provide a brief overview here. A more detailed discussion of related work can be found in Sec.A.
+2.1 Hallucination in LVLMs
+Hallucinations in LVLMs primarily concern cross-modal consistency Li et al.
+(2023c); Wang et al.
+(2023), which requires that the generated content remains faithful to the visual input.
+Among the various causes of hallucination, language priors Zhibo et al.
+(2023); Wu et al.
+(2022) are among the most fundamental.
+These priors originate from the LLM’s pretraining corpus and guide generation based on learned language patterns, which can produce fluent yet visually inconsistent outputs.
+Several approaches Chen et al.
+(2025); Liu et al.
+(2024a); Gunjal et al.
+(2024) attempt to mitigate language priors by constructing specialized training datasets, but these methods often incur significant additional training costs.
+Alternatively, recent training-free methods have explored contrastive decoding Li et al.
+(2023a); Chuang et al.
+(2024), typically constructing negative visual contexts based on the original image Leng et al.
+(2024); Wang et al.
+(2024c); Zhu et al.
+(2024b); Zhang et al.
+(2025) in an attempt to elicit and suppress language priors.
+In this work, we identify critical limitations of these approaches that stem from their reliance on modifying the original image.
+Specifically, such modifications may lead to visual information loss, biased language priors, and distribution mismatch.
+We further conduct an in-depth analysis of the cross-image consistency of language priors and, based on this insight, propose CICD.
+Additionally, we introduce a method for distinguishing between essential priors and detrimental priors, enabling fine-grained and precise mitigation of language priors in LVLMs.
+3 Preliminary
+Figure 4: We input the target image and a distinct image into the LVLM separately.
+During the step-by-step generation of the response, we compute the Jensen-Shannon (JS) divergence between the two logit distributions at each step to distinguish essential priors from detrimental ones.
+For detrimental priors, we apply contrastive decoding to suppress harmful language biases, while regular decoding is used for essential priors.
+3.1 Contrastive Decoding
+LVLMs take as input a concatenation of vision tokens from a vision encoder and an encoded text prompt.
+After forward propagation, the model projects the output into a logit distribution to compute conditional probabilities and generate responses in an autoregressive manner:
+$y_{t}\sim p_{\theta}(y_{t}\mid v,x,y_{<t})$ (1)
+$=$ $\ \mathrm{softmax}\left[\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})\right]$
+Here, $v$ and $x$ represent the visual and textual inputs, respectively. $\mathrm{logit}_{\theta}$ represents the logit distribution formulated by the LVLM with parameters $\theta$.
+Since LVLMs adopt pretrained LLMs as their backbone, they inherently inherit language priors that are independent of the visual input.
+As a result, tokens ranked higher in the output distribution $\mathrm{logit}_{\theta}$ may appear semantically plausible but can be objectively inconsistent with the visual evidence.
+This constitutes a fundamental cause of hallucinations.
+A commonly used approach to mitigate such priors is contrastive decoding Li et al.
+(2023a); Leng et al.
+(2024), which modifies the original image to construct a native visual context that elicits language priors.
+These priors are then suppressed by subtracting the logit distribution derived from the native context from that of the original context.
+$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,x,y_{<t})$ (2)
+$=$ $(1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})-\alpha\,\mathrm{logit}_{N}$
+Here, $\mathrm{logit}_{N}$ denotes the logit distribution obtained from the native context, and $\mathrm{logit}_{\mathrm{CD}}$ represents the final distribution used for next-token prediction.
+3.2 Formal Analysis
+The core component of contrastive decoding lies in the construction of the native context.
+Our preliminary investigation reveals that existing methods, which modify the original image, introduce several limitations (see Sec.1).
+Further observations and analysis show that language priors are inherent characteristics of LVLMs and remain consistent across different images, as illustrated in Fig.1.
+This insight leads to a natural alternative, where the native context is constructed using an image distinct from the original.
+Under this design, Eq.
+2 is refined as:
+$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,x,y_{<t})$ (3)
+$=$ $\ (1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$
+$\quad-\alpha\,\mathrm{logit}_{\theta}(y_{t}\mid v^{\prime},x,y_{<t})$
+where $v\textquoteright$ ’ is a different image from $v$. We further provide a formal analysis of the effectiveness of this approach.
+Most existing methods employ sampling decoding, where a token is randomly selected according to the logit distribution.
+We define the set of plausible tokens under a given logit distribution as $S=V\cup T$, where $V$ represents tokens related to visual content, and $T$ represents tokens unrelated to the image, i.e., those driven by language priors.
+The negative context constructed by a different image yields a new logit distribution, whose plausible token set is $S\textquoteright=V\textquoteright\cup T$, given that language priors remain unchanged across images.
+The contrastive decoding operation can thus be interpreted as subtracting one plausible token set from another:
+$S-S^{\prime}=(V\cup T)-(V^{\prime}\cup T)=V\mbox{ if }V^{\prime}\cap V=\emptyset$ (4)
+In other words, this subtraction removes all language priors that are unrelated to the visual input.
+This approach achieves three objectives simultaneously: (1) retention of all essential visual information, (2) precise elimination of language priors, and (3) compatibility with the input distribution expected by LVLMs.
+4 Method
 Leveraging the cross-image consistency of language priors, we propose CICD (Cross-Images Contrastive Decoding), which effectively mitigates language priors while preventing visual information loss.
@@
 4, CICD comprises three key steps: (1) retrieving a distinct image, (2) identifying detrimental language priors, and (3) removing these priors through contrastive decoding.
-Figure 4: The input image and another distinct image are fed into LVLM respectively with identical text-prefix.
-We calculate the JS discrepancy between the two logit distributions to discriminate essential priors and detrimental priors.
-For detrimental priors, we perform contrastive decoding to alleviate language priors, while applying regular decoding for essential priors.
-The final output of this case is shown in the right part.
-IV-A Image Retrieval
-To prevent visual information loss during contrastive decoding, we select an image that has limited information overlap with the input image.
-Since visual features are extracted by the LVLM’s vision encoder, an effective strategy is to measure the similarity between two images using the vision encoder to estimate their information overlap.
-In other words, the image with the lowest similarity to the input image is selected as the reference image.
-For better generalizability, we adopt the pre-trained CLIP [29] as the retriever.
-We vectorize all images in the dataset in advance, and during inference, retrieve the one with the minimum cosine similarity to the input image.
-Although vector retrieval is time-efficient given the computational cost of LVLMs, practical deployment still requires loading an additional retriever and maintaining a vector database.
-As illustrated in Fig.1, we observe that language priors primarily manifest in content words such as objects, attributes, and relations.
-These visual elements tend to exhibit minimal information overlap across natural images.
-As a result, randomly selecting an image can be an effective and lightweight alternative in practical implementations.
-We further discuss this in Sec.V-G.
-To ensure reproducibility, we adopt the retrieval-based approach for deterministic image selection throughout our experiments.
-IV-B Identifying Detrimental Priors
-Although language priors can lead LVLMs to generate hallucinated content, they are still essential for producing fluent and coherent text.
-Therefore, we categorize language priors into two types: essential priors and detrimental priors, where only the latter need to be eliminated.
-We define essential priors as purely linguistic attributes that are entirely independent of visual information.
-Specifically, for essential priors, the logit distributions produced by LVLMs should remain highly consistent across diverse visual inputs.
-This intrinsic property enables a clear distinction between essential and detrimental priors.
+4.1 Image Retrieval
+As discussed earlier, it is essential to select an image with limited information that overlaps with the input image.
+Since visual features are extracted by the vision encoder, a natural strategy is to measure the similarity between two images using the encoder itself to estimate their overlap.
+In other words, the image with the lowest similarity to the input image is selected as the reference.
+To ensure better generalizability, we adopt the pretrained CLIP model Radford et al.
+(2021) as the retriever.
+All images in the dataset are vectorized in advance, and during inference, the image with the lowest cosine similarity to the input is retrieved as the contrastive image.
+Although vector retrieval is relatively efficient given the computational cost of LVLMs, practical deployment still requires loading an additional retriever and maintaining a vector database.
+As illustrated in Fig.1, we observe that language priors primarily manifest in content words such as objects, attributes, and relationships.
+These visual elements tend to exhibit minimal overlap across natural images.
+Therefore, randomly selecting an image can serve as an effective and lightweight alternative for practical applications.
+We provide further discussion in Sec.C.2.
+Note that, to ensure reproducibility, we adopt the retrieval-based strategy for consistent and deterministic image selection throughout all experiments.
+4.2 Identifying Detrimental Priors
+Although language priors may cause LVLMs to generate hallucinated content, they remain essential for producing fluent and coherent text.
+To address this, we categorize language priors into two types: essential priors and detrimental priors, with only the latter requiring elimination.
+Essential priors are defined as purely linguistic patterns that are entirely independent of visual input.
+As discussed in Sec.
+1, the logit distributions associated with essential priors should remain highly consistent across diverse visual inputs.
+This intrinsic property provides a reliable basis for distinguishing between essential and detrimental priors.
 Figure 5: 500 random images from the MSCOCO dataset captioned by LLaVA-1.5.
 For each next-token prediction, we statistic the JS divergence (JSD) between two logit distributions produced by different images.
-We employ Jensen-Shannon (JS) divergence to measure the consistency between two logit distributions produced by different images:
+We employ the Jensen–Shannon (JS) divergence to measure the consistency between two logit distributions produced from different images:
 $\mathcal{D}_{\mathrm{JS}}=\mathrm{JS}\left(p_{\theta}(y_{t}\mid v,x,y_{<t})||p_{\theta}(y_{t}\mid v\textquoteright,x,y_{<t})\right)$ (5)
@@
 The statistical results are visualized in Fig.5.
-We observe that the JSD values can be clearly divided into two distinct ranges: $\log_{10}(\mathcal{D}{\mathrm{JS}})>-4$ and $\log_{10}(\mathcal{D}_{\mathrm{JS}})\leq-4$.
-Larger JSD values correspond to detrimental priors, while smaller values indicate essential priors.
+We observe that the JSD values can be clearly separated into two distinct ranges: $\log_{10}(\mathcal{D}{\mathrm{JS}})>-4$ and $\log_{10}(\mathcal{D}_{\mathrm{JS}})\leq-4$.
+Higher JSD values are associated with detrimental priors, while lower values correspond to essential priors, as shown in Fig.2 and Fig.3.
 Based on this observation, we set a threshold of $\gamma=-4$ to distinguish between the two types of priors.
-IV-C Cross-Images Contrastive Decoding
-The model performs regular decoding when essential priors are detected, and applies contrastive decoding to eliminate detrimental priors.
-For detrimental priors, we hypothesize that smaller distributional differences indicate stronger detrimental effects.
-Consequently, we define the contrastive coefficient $\alpha$ in Eq.
-3 as:
-$\alpha=1-\log_{10}(\mathcal{D}_{\mathrm{JS}})$ (6)
-Based on the observed density of $\mathcal{D}_{\mathrm{JS}}$ values, we clip $\alpha$ to the interval $[1,3]$ to ensure numerical stability.
-The final logit distribution is defined as follows:
-$\mathrm{logit}_{CICD}(y_{t}\mid v,v^{\prime},x,y_{<t})=$ (7)
-$\begin{cases}\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}\leq\gamma\\ \mathrm{logit}_{CD}(y_{t}\mid v,v^{\prime},x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}>\gamma\end{cases}$
-where $\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq. 1, and $\mathrm{logit}_{CD}(y_{t}\mid v,v^{\prime},x,y_{<t})$ is defined in Eq. 3.
-V Experiment
-V-A Benchmarks
-To evaluate the effectiveness of our method in eliminating language priors, we employ four multimodal hallucination benchmarks: two generative benchmarks (CHAIR [10] and DetailCaps [11]), one discriminative benchmark (POPE [14]), and one hybrid benchmark (AMBER [30]).
-CHAIR assesses the proportion of hallucinated objects, which are generated by the model but not present in the reference annotations.
-Following prior works, we randomly select 500 images from the MSCOCO [46] dataset as the test set.
-We additionally select another 500 images to construct a validation set for hyperparameter tuning.
-DetailCaps is a fine-grained image captioning benchmark, accompanied by ground-truth detail captions generated by GPT-4V [47], Gemini-1.5-Pro [48], and GPT-4o [49] for evaluation.
-It contains 4,870 images from multiple datasets, and we use a subset of 700 images from MSCOCO to conduct our experiments.
-POPE is a widely adopted benchmark for evaluating object hallucinations by prompting LVLMs to identify whether a specific object is present in the image.
-It consists of 1,500 images, including 1,000 from MSCOCO and 500 from GQA [50].
-AMBER combines generative and discriminative tasks, and is evaluated on a curated set of 1,004 images.
-In addition to image captioning, it includes 14,216 questions designed to assess hallucinations in object, attribute, and relation recognition.
-V-B Evaluated LVLMs
-We evaluate our method on six widely-used LVLMs: InstructBLIP [6], LLaVA-1.5 [31], LLaVA-Next [32], Qwen-VL-Chat [7], Qwen2-VL [33], and Qwen2.5-VL [8].
-InstructBLIP and LLaVA-1.5 are based on Vicuna-7B [3] as their language backbone, while LLaVA-Next is built upon LLaMA3.
-The Qwen-VL series models, including Qwen2-VL and Qwen2.5-VL, are developed based on the QwenLM framework.
-Regarding vision encoders, InstructBLIP, LLaVA-1.5, LLaVA-Next, and Qwen-VL-Chat all use the pretrained CLIP model [29], whereas Qwen2-VL and Qwen2.5-VL employ a fine-tuned ViT [35].
-All models are implemented at the 8B parameter scale.
-TABLE I: Results on POPE.
-Acc.
-and F1 represent the average accuracy and F1-score across three datasets (MSCOCO, A-OKVQA, and GQA), respectively.
-The bolded results denote the best in each setting and the underlined results denote the second-best.
 Setting Method LLaVA-1.5 InstructBLIP Qwen-VL-Chat
@@
 CICD 78.60 80.30 75.55 78.63 81.94 81.37
-V-C Baselines
-We compare our method with several training-free baselines that specifically target language priors.
-VCD [21] introduces noise into the input image to suppress language priors, and then applies contrastive decoding for their elimination.
-ICD [22] constructs negative logits using adversarial instructions and enhances positive signals through contrastive decoding.
-IBD [23] strengthens the model’s attention to visual features, thereby increasing the influence of image information during decoding.
-DeFG [26] converts the generated caption into an image and re-generates the caption from that image, leveraging the preserved visual content to reduce hallucinations.
-V-D Implementation Details
-We use sampling-based decoding for next-token prediction with default settings.
-Following [21], we incorporate adaptive plausibility constraints into the decoding process.
-We use the MSCOCO 2014 validation set as the image retrieval corpus, which contains 40,504 images.
-For retrieval, we adopt CLIP111openai/clip-vit-large-patch14-336 as the feature extractor and perform similarity-based retrieval using precomputed image embeddings.
-All experiments are conducted on a single 24GB NVIDIA RTX 3090 GPU.
-V-E Main Results
-Results on POPE
-POPE is a discriminative benchmark that incorporates three distinct datasets: MSCOCO, A-OKVQA, and GQA.
-Each dataset uses three different negative sampling settings: Random, Popular, and Adversarial.
-We conducted a comprehensive comparison between our method and other baselines across three distinct LVLMs, as detailed in Tab.
-I.
-For streamlined presentation, we report the average performance across all three datasets as the final result.
-Our method demonstrates superior overall performance, particularly in the most critical metric, achieving the highest F1 scores across all experimental settings.
-This suggests that we can utilize cross-image consistency to effectively alleviate language priors and promote LVLMs to convey faithful visual information.
-Additionally, our method demonstrates consistently superior performance across diverse LVLMs, evidencing remarkable generalization capability.
-As a training-free approach, it offers immediate deployability in practical applications.
-Another key finding reveals that our method achieves significantly enhanced performance in the Popular setting, demonstrating an average F1 score improvement of 2.84 over the second-best approach.
-The Popular setting uses the most frequent objects to construct negative samples, which makes LVLMs more susceptible to language priors.
-The most substantial performance improvements achieved by our method under this setting further demonstrate its exceptional efficacy in mitigating language priors.
-TABLE II: Results on the test set of CHAIR.
-MaxLen=512 indicates setting the maximum generation length of 512 tokens for the model.
-The bolded results denote the best in each setting and the underlined results denote the second-best.
-Setting Method LLaVA-1.5 InstructBLIP
+Table 1: Results on POPE.
+Acc.
+and F1 represent the average accuracy and F1-score across three datasets (MSCOCO, A-OKVQA, and GQA), respectively.
+The bolded results denote the best in each setting, and the underlined results denote the second-best.
+4.3 Cross-Images Contrastive Decoding
+After completing the above preparation, we integrate the proposed method into the contrastive decoding process.
+Specifically, the model performs regular decoding when essential priors are detected and applies contrastive decoding to suppress detrimental priors.
+In the case of detrimental priors, smaller distributional differences imply a stronger detrimental effect.
+Accordingly, we define the contrastive coefficient $\alpha$ in Eq.
+3 as: $\alpha=1-\log_{10}(\mathcal{D}_{\mathrm{JS}})$.
+Based on the observed density of $\mathcal{D}_{\mathrm{JS}}$ values, we clip $\alpha$ to the interval $[1,3]$ to ensure numerical stability.
+The final logit distribution is defined as follows:
+$\mathrm{logit}_{CICD}(y_{t}\mid v,x,y_{<t})=$ (6)
+$\begin{cases}\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}\leq\gamma\\ \mathrm{logit}_{CD}(y_{t}\mid v,x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}>\gamma\end{cases}$
+where $\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq. 1, and $\mathrm{logit}_{CD}(y_{t}\mid v,x,y_{<t})$ is defined in Eq. 3.
+5 Experiment
+MaxLen Method LLaVA-1.5 InstructBLIP
 […]
-MaxLen=512 Regular 54.6 16.4 72.6 106.8 62.6 19.5 66.9 113.6
+64 Regular 24.4 8.9 56.6 53.7 35.6 13.2 56.4 54.8
+VCD 25.0 8.3 59 54.0 32.2 10.3 60.6 55.2
+ICD 23.2 8.1 58.4 54.4 29.8 9.8 60.6 55.7
+IBD 21.2 6.9 58.8 54.5 27.8 9.2 60.3 56.1
+DeFG 22.4 7.2 58.2 54.4 32.4 11 59.4 56.0
+CICD 18.0 6.1 59.6 53.3 23.8 7.7 62.2 54.8
+512 Regular 54.6 16.4 72.6 106.8 62.6 19.5 66.9 113.6
 VCD 59.8 17.8 75.6 104 64.8 18.8 71.9 108.6
@@
 CICD 43.8 11.7 75.0 103.3 49.8 13.7 70.3 98.8
-MaxLen=64 Regular 24.4 8.9 56.6 53.7 35.6 13.2 56.4 54.8
-VCD 25.0 8.3 59 54.0 32.2 10.3 60.6 55.2
-ICD 23.2 8.1 58.4 54.4 29.8 9.8 60.6 55.7
-IBD 21.2 6.9 58.8 54.5 27.8 9.2 60.3 56.1
-DeFG 22.4 7.2 58.2 54.4 32.4 11 59.4 56.0
-CICD 18.0 6.1 59.6 53.3 23.8 7.7 62.2 54.8
-TABLE III: Results on AMBER.
-We set the maximun generation length to 512 for the models in this benchmark.
-AMBER metirc is calculated by $(100-CHAIR+F1)/2$.
-The bolded results denote the best in each setting and the underlined results denote the second-best.
+Table 2: Results on the test set of CHAIR. MaxLen denotes the maximum generation length.
+5.1 Experiment Settings
+Due to space constraints, we present only the key aspects of our experimental setup here. Detailed settings can be found in Sec.B.
+Benchmarks
+We evaluate the performance of CICD across four widely adopted multimodal hallucination benchmarks to validate its effectiveness on different tasks.
+These include two generative benchmarks (CHAIR Rohrbach et al.
+(2018) and DetailCaps Ye et al.
+(2025)), one discriminative benchmark (POPE Li et al.
+(2023b)), and one hybrid benchmark (AMBER Wang et al.
+(2024a)).
+Evaluated LVLMs
+We evaluate the generalizability of CICD across six LVLMs from three model families: InstructBLIP Dai et al.
+(2023), the LLaVA series (LLaVA-1.5 Liu et al.
+(2024c) and LLaVA-Next Liu et al.
+(2024d)), and the Qwen series (Qwen-VL-Chat Bai et al.
+(2023b), Qwen2-VL-Instruct Wang et al.
+(2024b), and Qwen2.5-VL-Instruct Bai et al.
+(2025)).
+Unless stated otherwise, all models are implemented at the 7B scale.
+Baselines
+We compare CICD with four recent training-free methods that specifically target language priors: VCD Leng et al.
+(2024), ICD Wang et al.
+(2024c), IBD Zhu et al.
+(2024b), and DeFG Zhang et al.
+(2025).
+All of these methods construct negative visual contexts based on the original image.
+In contrast, we identify several limitations of this approach and propose CICD as a more effective alternative.
+5.2 Implementation Details
+We use sampling decoding for next-token prediction with default settings.
+Following Leng et al.
+(2024), we incorporate adaptive plausibility constraints into the decoding process.
+We use the MSCOCO 2014 validation set as the image retrieval corpus, which contains 40,504 images.
+All experiments are conducted on an NVIDIA RTX 3090 GPU.
 Model Method Generative Discriminative
@@
 ICD 9.8 53.9 46.7 5.1 69.9 80.3 70.9 75.3 82.75
-IBD 9.0 56.1 45.1 4.6 54.2 86.3 34.2 49.00 70.00
+IBD 9.0 56.1 45.1 4.6 54.2 86.3 34.2 49.0 70.00
 DeFG 9.7 54.1 44.5 5.2 72.1 80.7 75.1 77.8 84.05
 CICD 7.1 53.6 35.0 2.3 72.8 85.4 70.1 77.0 84.95
+Table 3: Results on AMBER. The maximum generation length is 512. AMBER metirc is calculated by $(100-CHAIR+F1)/2$.
+5.3 Main Results
+Method LLaVA-1.5 InstructBLIP Qwen-VL-Chat
+[…]
+Regular 55.7 17.4 52.60 58.6 18 52.99 9.9 7.3 30.73
+VCD 55.7 16.8 52.91 59.6 18.9 53.20 7.3 5.9 31.28
+ICD 53.9 16.6 52.82 55.7 16.7 53.24 5.6 7 29.70
+IBD 54.1 15.8 52.48 56.4 15.5 54.14 4.6 2.9 29.31
+DeFG 55.7 16.6 52.72 59.1 17.4 53.06 5.6 3.9 30.21
+CICD 45.6 13.1 55.80 45.7 13.2 54.20 5.9 3.5 31.34
+Table 4: Results on DetailCaps. The maximum generation length is 512.
+Results on POPE
+As shown in Tab.
+1, our method achieves the highest F1 scores across all experimental settings.
+This suggests that we can utilize cross-image consistency to effectively alleviate language priors and promote LVLMs to convey faithful visual information.
+The Popular setting uses frequent objects for negative samples, increasing susceptibility to language priors.
+The strong performance gains of our method in this setting highlight its effectiveness in mitigating such priors.
+Model Method CHAIR(MaxLen=512) POPE (F1$\uparrow$)
+[…]
+LLaVA-Next Regular 41.8 11.9 60.2 85.47 83.14 79.87
+CICD 37.6 10.4 62.7 88.22 86.29 82.76
+Qwen2-VL-Instruct Regular 53.8 10.0 69.4 88.33 85.82 82.43
+CICD 50.6 9.4 73.6 90.67 88.07 84.53
+Qwen2.5-VL-Instruct Regular 41.2 11.5 65.7 87.83 85.66 83.12
+CICD 39.4 9.6 68.1 89.95 88.00 84.75
+Table 5: Results on more LVLMs. F1 is the average F1-score across three datasets (MSCOCO, A-OKVQA, and GQA). The bolded results denote the best in each setting.
 Results on CHAIR
-CHAIR is specifically designed to assess object hallucination in the image captioning task, which is one of the most language-prior-susceptible tasks in multimodal research.
-We evaluate our method in two settings with different maximum generation lengths; the results are shown in Tab.
-II.
-Our method demonstrates superior performance across all experimental conditions, with particularly pronounced improvements observed at longer generation lengths.
+The results in Tab.
+2 demonstrate our method’s superior performance across all experimental conditions, with particularly pronounced improvements observed at longer generation lengths.
 Longer context lengths amplify the model’s susceptibility to language priors, making these experimental results particularly compelling evidence of our method’s efficacy in mitigating such priors.
-Compared to baselines, our method’s principal advantage lies in its complete preservation of visual information, which is directly reflected in the superior experimental results.
-Additionally, we preserve the essential language priors, which are also important in generative tasks.
+Compared to baselines, our method’s principal advantage lies in its complete preservation of visual information and essential language priors, which is directly reflected in the superior performance in image captioning.
 Results on AMBER
-AMBER comprises generative tasks (image captioning) and discriminative tasks (image understanding).
-The results are shown in Tab.
-III.
-We can observe that our method achieves the lowest hallucination frequency in generative tasks while simultaneously attaining state-of-the-art performance on the comprehensive AMBER metric.
-This superior performance stems from our focused approach to language prior mitigation.
-This problem is particularly critical for generative tasks, where language priors manifest more severely than in discriminative tasks, explaining our method’s pronounced advantages in image captioning.
-TABLE IV: Results on DetailCaps. The bolded results denote the best in each setting and the underlined results denote the second-best.
-[…]
-LLaVA-1.5 InstructBLIP Qwen-VL-Chat
-Regular 52.60 52.99 30.73
-VCD 52.91 53.20 31.28
-ICD 52.82 53.24 29.70
-IBD 52.48 54.14 29.31
-DeFG 52.72 53.06 30.21
-CICD 55.80 54.20 31.34
+As illustrated in Tab.
+3, our method achieves the lowest hallucination frequency in generative tasks while simultaneously attaining state-of-the-art performance on the comprehensive AMBER metric.
+Language prior is particularly critical for generative tasks, where language priors manifest more severely than in discriminative tasks, explaining our method’s pronounced advantages in image captioning.
 Results on DetailCaps
-We further validate our method’s effectiveness on the image captioning task, which is more susceptible to language priors.
 Tab.
-IV presents the evaluation results on DetailCaps.
+4 presents the evaluation results on DetailCaps.
 The CAPTURE metric quantitatively evaluates caption quality through both lexical matching and semantic alignment, measuring the consistency of key content (entities, attributes, and relations) between generated captions and multiple ground-truth references.
 Our method achieves the best performance across all three evaluated LVLMs, demonstrating its robust capability to mitigate language priors and significantly enhance image captioning performance.
-TABLE V: Results on more LVLMs. F1 is the average F1-score across three datasets (MSCOCO, A-OKVQA, and GQA). The bolded results denote the best in each setting.
+5.4 Effectiveness on More LVLMs
+To comprehensively demonstrate the effectiveness and generalizability of our method, we select three of the most recent open-source LVLMs to conduct experiments.
+The results are presented in Tab.
+5.
+Our method demonstrates significant performance improvements in all experiments, consistently reducing hallucinations across both generative and discriminative tasks through effective mitigation of language priors.
+We also test our method on these LVLMs by DetailCaps and the results are shown in Tab.
+7.
+The improvements observed across different LVLMs confirm our method’s practical advantages: an elegantly simple framework coupled with state-of-the-art effectiveness.
+Further experiments will be discussed in Appendix B.
+6 Conclusion
+In this paper, we address the issue of language priors in LVLMs.
+Through a comprehensive investigation, we reveal that language priors are intrinsic to the model and remain consistent across different images.
+Building on this observation, we propose Cross-Image Contrastive Decoding (CICD) — a training-free method that effectively mitigates language priors while preserving visual information.
+Extensive experiments across multiple benchmarks and LVLM architectures demonstrate the effectiveness of CICD, particularly in image captioning tasks where language priors are most prominent.
+The simplicity and efficiency of CICD offer strong practical potential, enabling seamless integration with other hallucination mitigation methods to further enhance LVLM performance.
+This integration forms a key direction for our future work.
+7 Limitation
+We observe that language priors remain consistent across different images.
+Based on this insight, we propose the Cross-Images Contrastive Decoding (CICD) method to eliminate language priors.
+Our method has two main limitations compared to regular decoding.
+First, it requires an additional image that differs sufficiently from the original input.
+Second, both the original and the additional images must undergo separate forward passes to obtain their respective logit distributions for contrastive decoding.
+Addtional Image
+While the additional image is necessary, there is no strict requirement on its degree of difference from the original.
+Hallucinations typically occur in content-rich elements such as objects, attributes, and relationships, which naturally exhibit substantial variability across real-world images.
+Furthermore, our experiments demonstrate that the effectiveness of hallucination mitigation is not strongly correlated with image similarity.
+The experimental results demonstrate that randomly selected images perform comparably to retrieved ones, highlighting the practical applicability of our method.
+This flexibility significantly lowers the barrier for real-world deployment.
+Twice Forward Propagation
+Although performing twice forward propagations doubles the theoretical computational cost, in practice, we merge the two inputs into a single batch and perform one forward propagation with a doubled batch size to achieve the same effect.
+Under this implementation, our method incurs only a 10% increase in inference time and a 5% increase in memory usage, demonstrating strong potential for practical deployment.
+Appendix A Detailed Related Work
+A.1 Large Vision-Language Models
+Typically, Large Vision-Language Models (LVLMs) Dai et al.
+(2023); Liu et al.
+(2024c); Bai et al.
+(2023b); Zhu et al.
+(2024a) consist of a visual encoder Radford et al.
+(2021); Dosovitskiy et al.
+(2021) and a Large Language Model (LLM) backbone Touvron et al.
+(2023); Chiang et al.
+(2023).
+The visual encoder extracts visual features, which are then aligned with text tokens through a fusion module to form the input context for the LLM.
+Through multi-stage training, including modality alignment and instruction tuning Liu et al.
+(2023), LVLMs acquire the ability to perform various multimodal tasks such as image captioning Ye et al.
+(2025), visual question answering Yu et al.
+(2024), and visual grounding Lu et al.
+(2024).
+Despite significant advancements, LVLMs still suffer from severe hallucination problems, often generating content that contradicts the visual input.
+A.2 Hallucination in LVLMs
+Originally, hallucinations Ji et al.
+(2023); Liu et al.
+(2024b) were defined as undesirable outputs or non-factual information.
+In the multimodal domain, the study of hallucinations primarily focuses on cross-modal consistency Li et al.
+(2023c); Wang et al.
+(2023), requiring that LVLM-generated content remains faithful to the visual inputs.
+Multiple factors contribute to hallucinations in LVLMs, including biased visual attention Zhu et al.
+(2025); Li et al.
+(2025), attention misallocation Huang et al.
+(2024); Yin et al.
+(2025); Kang et al.
+(2025), and language priors Leng et al.
+(2024); Zhu et al.
+(2024b); Wang et al.
+(2024c).
+Among these, language priors are the most direct cause, making it difficult for LVLMs to accurately express their understanding of visual information.
+A.3 Language Priors
+As LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge embedded in LLMs, which are referred to as language priors Zhibo et al.
+(2023); Wu et al.
+(2022).
+Language priors specifically manifest when LVLMs generate content that is linguistically plausible but inconsistent with visual evidence.
+Some works Chen et al.
+(2025); Liu et al.
+(2024a); Gunjal et al.
+(2024) meticulously design specialized training data to mitigate this issue.
+However, such training-based approaches exhibit limited generalizability and cannot be easily transferred to broader application scenarios.
+Recently, a series of training-free approaches have directly reduced language priors using contrastive decoding Li et al.
+(2023a); Chuang et al.
+(2024), achieving remarkable performance.
+These methods construct an alternative logit distribution on top of the original one through techniques such as masking the image Leng et al.
+(2024), perturbing the instruction Wang et al.
+(2024c), augmenting the vision input Zhu et al.
+(2024b), or performing cross-modal conversion Zhang et al.
+(2025).
+During decoding, the two logit distributions are contrasted to eliminate language priors.
+While these methods effectively mitigate linguistic priors, they inevitably incur visual information loss during contrastive decoding due to information overlap, where both distributions share partial visual features.
+Through a comprehensive analysis of language priors, we identify their vision-independent nature.
+Based on this insight, we propose the Cross-Images Contrastive Decoding (CICD) method, which eliminates language priors by leveraging their consistency across different images, while preserving visual information through inter-image variation.
+Appendix B Detailed Experiment Settings
+B.1 Benchmarks
+In order to validate the effectiveness of our method in mitigating language hallucination, we evaluate its performance on four widely-used multimodal hallucination benchmarks: two generative benchmarks (CHAIR Rohrbach et al.
+(2018) and DetailCaps Ye et al.
+(2025)), one discriminative benchmark (POPE Li et al.
+(2023b)), and one hybrid benchmark (AMBER Wang et al.
+(2024a)).
+CHAIR
+assesses the proportion of hallucinated objects, which are generated by the model but not present in the reference annotations.
+Following prior works, we randomly select 500 images from the MSCOCO Lin et al.
+(2014) dataset as the test set.
+We additionally select another 500 images to construct a validation set for hyperparameter tuning.
+DetailCaps
+is a fine-grained image captioning benchmark, accompanied by ground-truth detail captions generated by GPT-4V Yang et al.
+(2023), Gemini-1.5-Pro Team et al.
+(2024), and GPT-4o OpenAI et al.
+(2024) for evaluation.
+It contains 4,870 images from multiple datasets, and we use a subset of 700 images from MSCOCO to conduct our experiments.
+POPE
+is a widely adopted benchmark for evaluating object hallucinations by prompting LVLMs to identify whether a specific object is present in the image.
+POPE incorporates three distinct datasets: MSCOCO, A-OKVQA, and GQA.
+Each dataset uses three different negative sampling settings: Random, Popular, and Adversarial.
+Finally, it consists of 9,000 questions and 1,500 images.
+AMBER
+combines generative and discriminative tasks, and is evaluated on a curated set of 1,004 images.
+In addition to image captioning, it includes 14,216 questions designed to assess hallucinations in object, attribute, and relation recognition.
+B.2 Evaluated LVLMs
+To demonstrate the generalizability of the proposed CICD as a broadly applicable plug-and-play module, we evaluate its performance across six different LVLMs spanning three model families: InstructBLIP Dai et al.
+(2023), LLaVA-1.5 Liu et al.
+(2024c), LLaVA-Next Liu et al.
+(2024d), Qwen-VL-Chat Bai et al.
+(2023b), Qwen2-VL-Instruct Wang et al.
+(2024b), and Qwen2.5-VL-Instruct Bai et al.
+(2025).
+InstructBLIP and LLaVA-1.5 are both built upon Vicuna-7B Chiang et al.
+(2023) as their language backbone, while LLaVA-Next is based on LLaMA3.
+The Qwen-VL series, including Qwen2-VL-Instruct and Qwen2.5-VL-Instruct, is developed using the QwenLM framework.
+As for vision encoders, InstructBLIP, LLaVA-1.5, LLaVA-Next, and Qwen-VL-Chat utilize the pretrained CLIP model Radford et al.
+(2021), whereas Qwen2-VL-Instruct and Qwen2.5-VL-Instruct adopt a fine-tuned ViT Dosovitskiy et al.
+(2021).
+All models are implemented at the 8B parameter scale.
+B.3 Baselines
+We compare our method with several training-free baselines that specifically target language priors.
+VCD Leng et al. (2024) introduces noise into the input image to suppress language priors, and then applies contrastive decoding for their elimination.
+ICD Wang et al. (2024c) constructs negative logits using adversarial instructions and enhances positive signals through contrastive decoding.
+IBD Zhu et al. (2024b) strengthens the model’s attention to visual features, thereby increasing the influence of image information during decoding.
+DeFG Zhang et al. (2025) converts the generated caption into an image and re-generates the caption from that image, leveraging the preserved visual content to reduce hallucinations.
+Appendix C More Experiment Results
+C.1 Effectiveness on Larger LVLMs
+We evaluate the effectiveness of our method on larger models using two benchmarks, POPE and CHAIR, with the results presented in Tab.6.
+Our method continues to deliver significant performance improvements on a 13B model, indicating its strong generalizability.
+These results suggest that our approach scales well and remains effective across different model sizes.
 Model Method CHAIR(MaxLen=512) POPE (F1$\uparrow$)
 […]
-LLaVA-next Regular 39.0 10.8 62.1 85.47 83.14 79.87
-CICD 37.6 10.4 62.7 88.22 86.29 82.76
-Qwen2-VL Regular 53.8 10.0 69.4 88.33 85.82 82.43
-CICD 35.0 9.2 60.5 90.67 88.07 84.53
-Qwen2.5-VL Regular 41.2 11.5 65.7 87.83 85.66 83.12
-CICD 37.2 9.3 60.0 89.95 88.00 84.75
-V-F Effectiveness on More LVLMs
-To comprehensively demonstrate the effectiveness and generalizability of our method, we select three of the most recent open-source LVLMs to conduct experiments.
-The results are presented in Tab.
-V.
-Our method demonstrates significant performance improvements in all experiments, consistently reducing hallucinations across both generative and discriminative tasks through effective mitigation of language priors.
-The improvements observed across different LVLMs confirm our method’s practical advantages: an elegantly simple framework coupled with state-of-the-art effectiveness.
-TABLE VI: This experiment conducted by LLaVA-1.5.
-Random Image and Retrieved Image represent the contrastive images selected through random sampling and image retrieval, respectively.
-w/o Image represents the contrastive distribution is constrcuted by removing the input image.
+LLaVA-1.5-13B Regular 58.8 17 73.4 82.94 81.46 78.49
+CICD 44.4 12.9 76.0 87.92 85.68 80.99
+INstructBLIP-13B Regular 65.6 21.0 62.7 81.31 77.20 74.69
+CICD 44 11.6 65.7 87.52 83.60 78.96
+Table 6: Results on larger LVLMs. The maximum generation length is 512.
+Method LLaVA-Next Qwen2-VL-Instruct Qwen2.5-VL-Instruct
+[…]
+Regular 61.06 38.0 10.4 60.04 51.3 10.6 62.62 42.2 11.7
+CICD 62.19 36.2 8.5 61.79 48.4 9.9 65.07 43.8 10.9
+Table 7: Results on DetailCaps with more LVLMs.
 […]
@@
 Retrieved Image 18.0 6.1 59.6
-V-G Divergence Between Images
+Table 8: This experiment was conducted by LLaVA-1.5.
+Random Image and Retrieved Image represent the contrastive images selected through random sampling and image retrieval, respectively.
+w/o Image represents the contrastive distribution that is constructed by removing the input image.
+Figure 6: The correlation between similarity and performance.
+We conducted this experiment on the validation set of CHAIR.
+Sim-CLIP represents normalized image similarity, measured as the cosine distance between the features of two images extracted by CLIP.
+C.2 Divergence Between Images
 Contrastive decoding is a method that suppresses consistent features between two distributions while amplifying distinctive ones.
@@
 Compared to regular decoding methods, images with varying similarity levels all demonstrate significant improvements.
-As discussed in Sec.IV-A, the influence of language priors is primarily reflected in the LVLMs’ comprehension of objects, attributes, and relationships.
+As discussed in Sec.4.1, the influence of language priors is primarily reflected in the LVLMs’ comprehension of objects, attributes, and relationships.
 This information exhibits limited overlap across real-world images.
-We hypothesize that random image selection could achieve comparable effectiveness, with experimental results presented in Tab.VI.
+We hypothesize that random image selection could achieve comparable effectiveness, with experimental results presented in Tab.8.
 Experimental results demonstrate that random image selection can also significantly mitigate language priors in LVLMs.
@@
 In Tab.
-VI, we also attempted to construct language priors by removing visual inputs, but found this approach less effective compared to eliminating linguistic priors through cross-image consistency.
+8, we also attempted to construct language priors by removing visual inputs, but found this approach less effective compared to eliminating linguistic priors through cross-image consistency.
 We attribute this limitation to LVLMs being trained on large-scale multimodal data, where removing visual inputs disrupts the model’s intrinsic representations and leads to information distortion.
 Consequently, language priors in LVLMs cannot be directly constructed, which further substantiates the necessity of leveraging cross-image consistency in our method.
-Figure 6: The correlation between similarity and performance.
-We conduct this experiment on the validation set of CHAIR.
-Sim-CLIP represents normalized image similarity, measured as the cosine distance between the features of two images extracted by CLIP.
 Figure 7: The results on the validation set of CHAIR using LLaVA-1.5. The X-axis represents the value of $\gamma$.
-V-H Threshold of Detrimental Priors
+C.3 Threshold of Detrimental Priors
 While language priors contribute to hallucinations in LVLMs, they also play a critical role in maintaining textual coherence and fluency.
@@
 Another notable finding is that preserving essential priors proves meaningful, as it ensures fluent and coherent generation contexts.
-VI Conclusion
-In this paper, we focus on the language priors issue in LVLMs.
-Through a comprehensive investigation, we reveal that language priors are intrinsic characteristics of the model and remain consistent across different images.
-Based on this observation, we propose Cross-Images Contrastive Decoding (CICD)—a training-free method that effectively mitigates language priors while preserving visual information.
-Extensive experiments across multiple benchmarks and LVLM architectures validate the effectiveness of CICD, particularly in image captioning tasks where language priors are most pronounced.
-The simplicity and efficiency of CICD give it substantial practical potential, enabling synergistic integration with other hallucination mitigation methods to further enhance LVLM performance.
-This constitutes our primary direction for future work.

### v3

@@
-Mitigate Language Priors in Large Vision-Language Models by Cross-Images Contrastive Decoding
+Cross-Image Contrastive Decoding: Precise, Lossless Suppression of Language Priors in Large Vision-Language Models
 Abstract
 Language priors are a major cause of hallucinations in Large Vision-Language Models (LVLMs), often leading to text that is linguistically plausible but visually inconsistent.
-Recent work explores contrastive decoding as a training-free solution, but these methods typically construct negative visual contexts from the original image, resulting in visual information loss and distorted distribution.
-Motivated by the observation that language priors stem from the LLM backbone and remain consistent across images, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective training-free method that uses different images to construct negative visual contexts.
-We further analyze the cross-image behavior of language priors and introduce a distinction between essential priors (supporting fluency) and detrimental priors (causing hallucinations), enabling selective suppression.
+Recent work explores contrastive decoding as a training-free solution, but these methods typically construct negative contexts from the original image, resulting in visual information loss and distorted distribution.
+Motivated by the observation that language priors stem from the LLM backbone and remain consistent across images, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective training-free method that uses different images to construct negative contexts.
+We further analyze the cross-image behavior of language priors and introduce a distinction between essential priors (supporting fluency) and detrimental priors (causing hallucinations).
 By selectively preserving essential priors and suppressing detrimental ones, our method reduces hallucinations while maintaining coherent and fluent language generation.
-Experiments on four benchmarks and six LVLMs across three model families confirm the effectiveness and generalizability of CICD, especially in image captioning, where language priors are particularly pronounced.
-Code will be released upon acceptance.
-Mitigate Language Priors in Large Vision-Language Models by Cross-Images Contrastive Decoding
-Jianfei Zhao1,2, Feng Zhang1, Xin Sun1, Chong Feng1,3,thanks: Corresponding Autor: chongfeng@bit.edu.cn , 1School of Computer Science and Technology, Beijing Institute of Technology, Beijing, China, 2Zhongguancun Academy, Beijing, China, 3Southeast Academy of Information Technology, Beijing Institute of Technology, Fujian, China,
+Experiments on 4 benchmarks and 6 LVLMs across three model families confirm the effectiveness and generalizability of CICD, especially in image captioning, where language priors are particularly pronounced.
+Code will be released once accepted.
+Cross-Image Contrastive Decoding: Precise, Lossless Suppression of Language Priors in Large Vision-Language Models
+Jianfei Zhao1,2, Feng Zhang1, Xin Sun1, Chong Feng1,3,thanks: Corresponding Autor: fengchong@bit.edu.cn , 1School of Computer Science and Technology, Beijing Institute of Technology, Beijing, China, 2Zhongguancun Academy, Beijing, China, 3Southeast Academy of Information Technology, Beijing Institute of Technology, Fujian, China,
 1 Introduction
-Figure 1: A case of language prior. Adding the same prefix when captioning another image will leads to the same language priors. The model used is LLaVA-1.5.
+Figure 1: A case of language prior. Adding the same prefix when captioning another image will lead to the same language priors. The model used is LLaVA-1.5.
 Recent advances in Large Language Models (LLMs) Brown et al.
@@
 (2024b) being the most fundamental underlying cause.
-Since LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge encoded in the LLMs.
+Specifically, since LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge encoded in the LLMs.
 As a result, their outputs are influenced not only by visual evidence but also by pre-learned language patterns.
@@
 For instance, when describing an office scene, an LVLM may erroneously mention a "computer" even if no such object is present in the image.
+Figure 2: The overlap ratio of image-related words between responses generated from negative contexts and those from regular inputs.
+Both VCD and ICD exhibit over 30% overlap, indicating that their negative contexts retain a considerable amount of essential visual information.
+In contrast, our method results in a nearly negligible overlap.
+The detailed analysis methodology can be found in Appendix.
+D.
+Methods for mitigating language priors in LVLMs can generally be classified into two categories.
+The first involves constructing specialized training datasets to reduce the model’s dependence on linguistic biases Zhibo et al.
+(2023); Wu et al.
+(2022); Chen et al.
+(2025), though such approaches often incur substantial training costs.
+Another line of work explores plug-and-play, training-free methods, which are also the focus of our study.
+Among them, many recent studies adopt contrastive decoding Li et al.
+(2023a), which constructs negative contexts to reveal and suppress language priors.
+Some approaches directly perturb the input image Leng et al.
+(2024); Wang et al.
+(2024c) to alter the visual context, while others enhance the visual features extracted from the original image Zhu et al.
+(2024b); Zhang et al.
+(2025).
+These methods typically compute the difference between the logit distributions obtained from the original and perturbed (i.e., negative) contexts to reduce the model’s reliance on language priors.
+Overall, following the paradigm introduced in VCD Leng et al.
+(2024), these methods routinely perturb the original inputs to generate negative contexts.
+However, this prevailing design choice raises a critical question: Is perturbing the original context to construct negative contexts truly the most effective strategy for inducing and mitigating language priors?
 Figure 3: Essential priors (left) and detrimental priors (right).
@@
 (2021).
-Methods for mitigating language priors in LVLMs can generally be classified into two categories.
-The first involves constructing specialized training datasets to reduce the model’s dependence on linguistic biases Zhibo et al.
-(2023); Wu et al.
-(2022); Chen et al.
-(2025), though such approaches often incur substantial training costs.
-In this work, we focus on plug-and-play, training-free methods.
-Many recent studies adopt contrastive decoding Li et al.
-(2023a), which constructs negative contexts to expose and suppress language priors.
-Some approaches directly perturb the input image Leng et al.
-(2024); Wang et al.
-(2024c) to alter the visual context, while others enhance the visual features extracted from the original image Zhu et al.
-(2024b); Zhang et al.
-(2025).
-These methods typically compute the difference between the logit distributions obtained from the original and modified (i.e., negative) contexts to reduce the model’s reliance on language priors.
-Following the paradigm introduced in VCD Leng et al.
-(2024), these methods routinely modify the original image to generate negative visual contexts.
-However, this prevailing design choice raises a critical question: Is constructing negative contexts through perturbations of the original image truly the most effective strategy for inducing and mitigating language priors?
 We argue that the answer is unequivocally negative.
-Constructing negative contexts based on the original image can lead to two main issues: (1) Visual information loss, where the negative visual context retains partial visual information, which may be degraded during contrastive decoding.
+Constructing negative contexts based on the original inputs can lead to two main issues: (1) Visual information loss, where the negative context retains partial visual information, which may be degraded during contrastive decoding.
+As shown in Fig.
+2, compared to regular responses, those generated from negative contexts retain over 30% of the essential visual information, which is subsequently weakened during the contrastive decoding process.
 (2) Distorted distribution, where perturbing the model’s input to induce priors may disrupt its internal representations, ultimately resulting in a distorted logit distribution.
-This analysis raises an important question: How can the above issues be avoided when constructing negative contexts?
-In fact, simply taking another image to construct negative context is sufficient.
+This analysis raises an important question: How should negative contexts be constructed to avoid the above issue?
+Figure 4: A case of JS divergence between the logit distributions of two images.
+In fact, simply taking a different image to construct negative context is sufficient.
 This insight is based on the thought that the language priors guiding step-wise token generation in LVLMs originate from the LLM backbone and are largely independent of the input image.
 As shown in Fig.1, LVLM produces highly consistent outputs when given completely different images but the same prefix, which strongly supports our hypothesis.
-Building on this, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective method for mitigating language priors in LVLMs.
-Instead of modifying the original image, CICD constructs the negative context using a different image that shares the same language priors.
+Building on this, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective method for suppressing language priors in LVLMs.
+Instead of perturbing the original inputs, CICD constructs the negative context using a different image that shares the same language priors.
 Our method avoids the two key issues present in previous approaches.
 Since the overlap between the original and the new image is limited, our approach avoids introducing critical visual information into the negative context.
+As shown in Fig.
+2, only a negligible amount (approximately 0.11%) of essential visual information is retained in our negative context.
 Moreover, it enables the model to reveal intrinsic priors, maintaining the model’s natural inference process without injecting external bias.
@@
 3.2.
-Figure 3: A case of JS divergence between the logit distributions of two images.
 Focusing on language priors, we further explore whether all such priors should be strictly eliminated.
-We argue that the answer is negative.
+The answer is intuitively negative.
 Language priors can be divided into two categories: (1) Essential Priors, which are necessary for maintaining textual fluency, and (2) Detrimental Priors, which introduce information unrelated to the visual input.
@@
 Specifically, we compute the Jensen–Shannon (JS) divergence between the logit distributions of different visual contexts, and apply a stable threshold to identify essential priors that should be retained.
-We validate the effectiveness of CICD on four widely used benchmarks: CHAIR Rohrbach et al.
-(2018), POPE Li et al.
-(2023b), AMBER Wang et al.
-(2024a), and DetailCaps Ye et al.
-(2025).
-To assess the generalizability of our approach, we conduct experiments across six LVLMs from three different model families: InstructBLIP Dai et al.
-(2023), LLaVA-1.5 Liu et al.
-(2024c), LLaVA-Next Liu et al.
-(2024d), Qwen-VL-Chat Bai et al.
-(2023b), Qwen2-VL-Instruct Wang et al.
-(2024b), and Qwen2.5-VL-Instruct Bai et al.
-(2025).
-Experimental results show that CICD consistently reduces hallucinations by effectively mitigating language priors in LVLMs, outperforming state-of-the-art methods across all benchmarks and models.
+Extensive experiments on 4 widely used benchmarks and 6 LVLMs across three model families demonstrate the effectiveness and generalizability of our method.
+By suppressing language priors, CICD consistently reduces hallucinations and achieves state-of-the-art performance across all benchmarks.
 The improvement is particularly pronounced in the image captioning task, which tends to suffer more from language priors due to its longer textual outputs.
 In summary, our main contributions are as follows:
-  • We conduct an in-depth analysis of existing contrastive decoding methods and identify their inherent limitations rooted in the reliance on the original image.
-We further verify the consistency of language priors across images and propose a stable approach for distinguishing essential priors from detrimental ones.
+  • We analyze existing contrastive decoding methods in depth and reveal their limitations stemming from reliance on the original image.
+To address this, we verify the consistency of language priors across images and propose a stable approach to distinguish essential priors from detrimental ones.
   • Based on this analysis, we propose CICD (Cross-Images Contrastive Decoding), which leverages retrieved images to induce language priors while selectively preserving essential priors during the mitigation process.
@@
 2 Related Work
-Due to space limitations, we provide a brief overview here. A more detailed discussion of related work can be found in Sec.A.
+Due to space limitations, we provide a brief overview here. A more detailed discussion of related work can be found in Appendix.A.
 2.1 Hallucination in LVLMs
@@
 (2023); Wu et al.
-(2022) are among the most fundamental.
+(2022) are the most fundamental.
 These priors originate from the LLM’s pretraining corpus and guide generation based on learned language patterns, which can produce fluent yet visually inconsistent outputs.
@@
 (2023a); Chuang et al.
-(2024), typically constructing negative visual contexts based on the original image Leng et al.
+(2024), typically constructing negative contexts based on the original image Leng et al.
 (2024); Wang et al.
@@
 In this work, we identify critical limitations of these approaches that stem from their reliance on modifying the original image.
-Specifically, such modifications may lead to visual information loss, biased language priors, and distribution mismatch.
+Specifically, such modifications may lead to visual information loss and distorted distribution.
 We further conduct an in-depth analysis of the cross-image consistency of language priors and, based on this insight, propose CICD.
@@
 3 Preliminary
-Figure 4: We input the target image and a distinct image into the LVLM separately.
+Figure 5: The diagram of our method.
+We input the target image and a distinct image into the LVLM separately.
 During the step-by-step generation of the response, we compute the Jensen-Shannon (JS) divergence between the two logit distributions at each step to distinguish essential priors from detrimental ones.
@@
 $=$ $\ \mathrm{softmax}\left[\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})\right]$
-Here, $v$ and $x$ represent the visual and textual inputs, respectively. $\mathrm{logit}_{\theta}$ represents the logit distribution formulated by the LVLM with parameters $\theta$.
-Since LVLMs adopt pretrained LLMs as their backbone, they inherently inherit language priors that are independent of the visual input.
+Here, $v$ and $x$ represent the visual and textual inputs, respectively. $\mathrm{logit}_{\theta}$ represents the logit distribution formulated by the model with parameters $\theta$.
+Since LVLMs adopt pretrained LLMs as their backbone, they inherently inherit language priors independent of the visual input.
 As a result, tokens ranked higher in the output distribution $\mathrm{logit}_{\theta}$ may appear semantically plausible but can be objectively inconsistent with the visual evidence.
@@
 As illustrated in Fig.
-4, CICD comprises three key steps: (1) retrieving a distinct image, (2) identifying detrimental language priors, and (3) removing these priors through contrastive decoding.
+5, CICD comprises three key steps: (1) retrieving a distinct image, (2) identifying detrimental language priors, and (3) removing these priors through contrastive decoding.
 4.1 Image Retrieval
@@
 Therefore, randomly selecting an image can serve as an effective and lightweight alternative for practical applications.
-We provide further discussion in Sec.C.2.
+We provide further discussion in Appendix.C.2.
 Note that, to ensure reproducibility, we adopt the retrieval-based strategy for consistent and deterministic image selection throughout all experiments.
@@
 The statistical results are visualized in Fig.6.
-We observe that the JSD values can be clearly separated into two distinct ranges: $\log_{10}(\mathcal{D}{\mathrm{JS}})>-4$ and $\log_{10}(\mathcal{D}_{\mathrm{JS}})\leq-4$.
+We observe that the JSD values can be clearly separated into two distinct ranges: $\log_{10}(\mathcal{D}_{\mathrm{JS}})>-4$ and $\log_{10}(\mathcal{D}_{\mathrm{JS}})\leq-4$.
 Higher JSD values are associated with detrimental priors, while lower values correspond to essential priors, as shown in Fig.3 and Fig.4.
 Based on this observation, we set a threshold of $\gamma=-4$ to distinguish between the two types of priors.
+Further discussion can be found in Appendix.
+C.3
 Setting Method LLaVA-1.5 InstructBLIP Qwen-VL-Chat
@@
 CICD 43.8 11.7 75.0 103.3 49.8 13.7 70.3 98.8
-Table 2: Results on the test set of CHAIR. MaxLen denotes the maximum generation length.
-5.1 Experiment Settings
-Due to space constraints, we present only the key aspects of our experimental setup here. Detailed settings can be found in Sec.B.
-Benchmarks
-We evaluate the performance of CICD across four widely adopted multimodal hallucination benchmarks to validate its effectiveness on different tasks.
-These include two generative benchmarks (CHAIR Rohrbach et al.
-(2018) and DetailCaps Ye et al.
-(2025)), one discriminative benchmark (POPE Li et al.
-(2023b)), and one hybrid benchmark (AMBER Wang et al.
-(2024a)).
-Evaluated LVLMs
-We evaluate the generalizability of CICD across six LVLMs from three model families: InstructBLIP Dai et al.
-(2023), the LLaVA series (LLaVA-1.5 Liu et al.
-(2024c) and LLaVA-Next Liu et al.
-(2024d)), and the Qwen series (Qwen-VL-Chat Bai et al.
-(2023b), Qwen2-VL-Instruct Wang et al.
-(2024b), and Qwen2.5-VL-Instruct Bai et al.
-(2025)).
-Unless stated otherwise, all models are implemented at the 7B scale.
-Baselines
-We compare CICD with four recent training-free methods that specifically target language priors: VCD Leng et al.
-(2024), ICD Wang et al.
-(2024c), IBD Zhu et al.
-(2024b), and DeFG Zhang et al.
-(2025).
-All of these methods construct negative visual contexts based on the original image.
-In contrast, we identify several limitations of this approach and propose CICD as a more effective alternative.
-5.2 Implementation Details
-We use sampling decoding for next-token prediction with default settings.
-Following Leng et al.
-(2024), we incorporate adaptive plausibility constraints into the decoding process.
-We use the MSCOCO 2014 validation set as the image retrieval corpus, which contains 40,504 images.
-All experiments are conducted on an NVIDIA RTX 3090 GPU.
+Table 2: Results on the test set of CHAIR. MaxLen denotes the maximum generation length. Cs and Ci represent CHAIRs and CHAIRi respectively.
 Model Method Generative Discriminative
@@
 Table 3: Results on AMBER. The maximum generation length is 512. AMBER metirc is calculated by $(100-CHAIR+F1)/2$.
+5.1 Experiment Settings
+Due to space constraints, we present only the key aspects of our experimental setup here. Detailed settings can be found in Sec.B.
+Benchmarks
+We evaluate the performance of CICD across 4 widely adopted multimodal hallucination benchmarks to validate its effectiveness on different tasks.
+These include two generative benchmarks (CHAIR Rohrbach et al.
+(2018) and DetailCaps Ye et al.
+(2025)), one discriminative benchmark (POPE Li et al.
+(2023b)), and one hybrid benchmark (AMBER Wang et al.
+(2024a)).
+Evaluated LVLMs
+We evaluate the generalizability of CICD across 6 LVLMs from three model families: InstructBLIP Dai et al.
+(2023), the LLaVA series (LLaVA-1.5 Liu et al.
+(2024c) and LLaVA-Next Liu et al.
+(2024d)), and the Qwen series (Qwen-VL-Chat Bai et al.
+(2023b), Qwen2-VL-Instruct Wang et al.
+(2024b), and Qwen2.5-VL-Instruct Bai et al.
+(2025)).
+Unless stated otherwise, all models are implemented at the 7B scale.
+Baselines
+We compare CICD with 4 recent training-free methods that specifically target language priors: VCD Leng et al.
+(2024), ICD Wang et al.
+(2024c), IBD Zhu et al.
+(2024b), and DeFG Zhang et al.
+(2025).
+All of these methods construct negative contexts based on the original image.
+In contrast, we identify several limitations of this approach and propose CICD as a more effective alternative.
+5.2 Implementation Details
+We use sampling decoding for next-token prediction with default settings.
+Following Leng et al.
+(2024), we incorporate adaptive plausibility constraints into the decoding process.
+We use the MSCOCO 2014 validation set as the image retrieval corpus, which contains 40,504 images.
+All experiments are conducted on an NVIDIA RTX 3090 GPU.
 5.3 Main Results
@@
 […]
-Regular 55.7 17.4 52.60 58.6 18 52.99 9.9 7.3 30.73
+Regular 55.7 17.4 52.60 58.6 18.0 52.99 9.9 7.3 30.73
 VCD 55.7 16.8 52.91 59.6 18.9 53.20 7.3 5.9 31.28
-ICD 53.9 16.6 52.82 55.7 16.7 53.24 5.6 7 29.70
+ICD 53.9 16.6 52.82 55.7 16.7 53.24 5.6 7.0 29.70
 IBD 54.1 15.8 52.48 56.4 15.5 54.14 4.6 2.9 29.31
@@
 CICD 45.6 13.1 55.80 45.7 13.2 54.20 5.9 3.5 31.34
-Table 4: Results on DetailCaps. The maximum generation length is 512.
+Table 4: Results on the COCO subset of DetailCaps. The maximum generation length is 512. Cs, Ci, and CAP represent CHAIRs, CHAIRi, and CAPTURE respectively.
 Results on POPE
@@
 CICD 39.4 9.6 68.1 89.95 88.00 84.75
-Table 5: Results on more LVLMs. F1 is the average F1-score across three datasets (MSCOCO, A-OKVQA, and GQA). The bolded results denote the best in each setting.
+Table 5: Results on more LVLMs. F1 is the average F1-score across three datasets (MSCOCO, A-OKVQA, and GQA).
 Results on CHAIR
@@
 We also test our method on these LVLMs by DetailCaps and the results are shown in Tab.
-7.
+6.
 The improvements observed across different LVLMs confirm our method’s practical advantages: an elegantly simple framework coupled with state-of-the-art effectiveness.
-Further experiments will be discussed in Appendix B.
+Further experiments will be discussed in Sec. B.
 6 Conclusion
-In this paper, we address the issue of language priors in LVLMs.
+In this paper, we mitigate the influence of language priors in LVLMs.
 Through a comprehensive investigation, we reveal that language priors are intrinsic to the model and remain consistent across different images.
@@
 (2024a)).
+Method LLaVA-Next Qwen2-VL-Instruct Qwen2.5-VL-Instruct
+[…]
+Regular 61.06 38.0 10.4 60.04 51.3 10.6 62.62 42.2 11.7
+CICD 62.19 36.2 8.5 61.79 48.4 9.9 65.07 43.8 10.9
+Table 6: Results on the COCO subset of DetailCaps with more LVLMs.
 CHAIR
-assesses the proportion of hallucinated objects, which are generated by the model but not present in the reference annotations.
+evaluates the proportion of hallucinated objects, which are generated by the model but not present in the reference annotations.
 Following prior works, we randomly select 500 images from the MSCOCO Lin et al.
@@
 We additionally select another 500 images to construct a validation set for hyperparameter tuning.
+This benchmark includes two metrics: CHAIRs and CHAIRi, defined as follows:
+CHAIRs $=\frac{|\text{Hallucinated Objects}|}{|\text{All Objects}|},$ (7)
+CHAIRi $=\frac{|\text{Hallucinated Sentences}|}{|\text{All Sentences}|}$
 DetailCaps
@@
 (2024) for evaluation.
-It contains 4,870 images from multiple datasets, and we use a subset of 700 images from MSCOCO to conduct our experiments.
+It contains 4,870 images collected from multiple datasets, and we use a subset of 700 images from MSCOCO to conduct our experiments.
+This benchmark primarily adopts the CAPTURE metric to evaluate caption quality.
+CAPTURE evaluates the alignment between generated and reference captions by calculating F1 scores using both hard and soft matching across entities ($F1_{obj}$), attributes ($F1_{attr}$), and relations ($F1_{rel}$).
+The final score is computed through a weighted aggregation of these components:
+$CAPTURE=\frac{\alpha F1_{obj}+\beta F1_{attr}+\gamma F1_{rel}}{\alpha+\beta+\gamma}$ (8)
+where $\alpha=5,\beta=5,\gamma=2$
 POPE
 is a widely adopted benchmark for evaluating object hallucinations by prompting LVLMs to identify whether a specific object is present in the image.
-POPE incorporates three distinct datasets: MSCOCO, A-OKVQA, and GQA.
+It comprises three distinct datasets: MSCOCO, A-OKVQA, and GQA.
 Each dataset uses three different negative sampling settings: Random, Popular, and Adversarial.
-Finally, it consists of 9,000 questions and 1,500 images.
+In total, it includes 9,000 questions and 1,500 images.
+Accuracy and F1 score are used as the primary evaluation metrics.
 AMBER
@@
 In addition to image captioning, it includes 14,216 questions designed to assess hallucinations in object, attribute, and relation recognition.
+AMBER contians mulitple metrics: CHAIR, Cover, Hal, Cog.
+It provides an annotated objects list $A_{obj}={obj_{1}^{A},obj_{2}^{A},\cdots,obj_{n}^{A}}$, and the generated objects are labeled as $R^{\prime}_{obj}$.
+Each metric is calculated as follows:
+CHAIR $=1-\frac{\mathrm{len}(R^{\prime}_{obj}\cap A_{obj})}{\mathrm{len}(R^{\prime}_{obj})},$ (9)
+Cover $=\frac{\mathrm{len}(R^{\prime}_{obj}\cap A_{obj})}{\mathrm{len}(A_{obj})},$
+Hal =$=\frac{\{\text{CHAIR}>0\}}{\{\text{All Caps}\}},$
+Cog $=\frac{\mathrm{len}(R^{\prime}_{obj}\cap H_{obj})}{\mathrm{len}(R^{\prime}_{obj})},$
+where $H_{obj}$ denotes the set of hallucinated target objects generated by the LVLMs, and All Caps refers to all generated captions.
 B.2 Evaluated LVLMs
@@
 All models are implemented at the 8B parameter scale.
+Model Method CHAIR(MaxLen=512) POPE (F1$\uparrow$)
+[…]
+LLaVA-1.5-13B Regular 58.8 17.0 73.4 82.94 81.46 78.49
+CICD 44.4 12.9 76.0 87.92 85.68 80.99
+INstructBLIP-13B Regular 65.6 21.0 62.7 81.31 77.20 74.69
+CICD 44.0 11.6 65.7 87.52 83.60 78.96
+Table 7: Results on larger LVLMs. The maximum generation length is 512.
 B.3 Baselines
@@
 Appendix C More Experiment Results
+Figure 7: The correlation between similarity and performance.
+We conducted this experiment on the validation set of CHAIR.
+Sim-CLIP represents normalized image similarity, measured as the cosine distance between the features of two images extracted by CLIP.
 C.1 Effectiveness on Larger LVLMs
@@
 These results suggest that our approach scales well and remains effective across different model sizes.
-Model Method CHAIR(MaxLen=512) POPE (F1$\uparrow$)
-[…]
-LLaVA-1.5-13B Regular 58.8 17 73.4 82.94 81.46 78.49
-CICD 44.4 12.9 76.0 87.92 85.68 80.99
-INstructBLIP-13B Regular 65.6 21.0 62.7 81.31 77.20 74.69
-CICD 44 11.6 65.7 87.52 83.60 78.96
-Table 6: Results on larger LVLMs. The maximum generation length is 512.
-Method LLaVA-Next Qwen2-VL-Instruct Qwen2.5-VL-Instruct
-[…]
-Regular 61.06 38.0 10.4 60.04 51.3 10.6 62.62 42.2 11.7
-CICD 62.19 36.2 8.5 61.79 48.4 9.9 65.07 43.8 10.9
-Table 7: Results on DetailCaps with more LVLMs.
-[…]
-CHAIR(MaxLen=512)
-Regular 54.6 16.4 72.6
-w/o Image 54.6 15.5 76.6
-Random Image 46.6 13.2 76.7
-Retrieved Image 43.8 11.7 75
-CHAIR(MaxLen=64)
-Regular 24.4 8.9 56.6
-w/o Image 25.4 7.6 61.9
-Random Image 21.6 6.9 58.8
-Retrieved Image 18.0 6.1 59.6
-Table 8: This experiment was conducted by LLaVA-1.5.
-Random Image and Retrieved Image represent the contrastive images selected through random sampling and image retrieval, respectively.
-w/o Image represents the contrastive distribution that is constructed by removing the input image.
-Figure 6: The correlation between similarity and performance.
-We conducted this experiment on the validation set of CHAIR.
-Sim-CLIP represents normalized image similarity, measured as the cosine distance between the features of two images extracted by CLIP.
 C.2 Divergence Between Images
@@
 We conduct experiments to explore the correlation between similarity and performance, as shown in Fig.
-6.
+7.
 We observe that the similarity between images exhibits no strong correlation with performance.
@@
 Consequently, language priors in LVLMs cannot be directly constructed, which further substantiates the necessity of leveraging cross-image consistency in our method.
-Figure 7: The results on the validation set of CHAIR using LLaVA-1.5. The X-axis represents the value of $\gamma$.
+[…]
+CHAIR(MaxLen=512)
+Regular 54.6 16.4 72.6
+w/o Image 54.6 15.5 76.6
+Random Image 46.6 13.2 76.7
+Retrieved Image 43.8 11.7 75.0
+CHAIR(MaxLen=64)
+Regular 24.4 8.9 56.6
+w/o Image 25.4 7.6 61.9
+Random Image 21.6 6.9 58.8
+Retrieved Image 18.0 6.1 59.6
+Table 8: This experiment was conducted by LLaVA-1.5.
+Random Image and Retrieved Image represent the contrastive images selected through random sampling and image retrieval, respectively.
+w/o Image represents the contrastive distribution that is constructed by removing the input image.
 C.3 Threshold of Detrimental Priors
@@
 A comprehensive analysis is shown in Fig.
-7.
+8.
 The experimental results align with the statistical analysis, demonstrating optimal performance at $\gamma=-4$.
 Another notable finding is that preserving essential priors proves meaningful, as it ensures fluent and coherent generation contexts.
+Appendix D Analysis of Visual Information Overlap
+Fig. 2 presents the overlap ratio of image-related words between responses generated from negative contexts and those from regular decoding. Below, we detail the analysis procedure.
+First, we apply several contrastive decoding methods, each constructing its own negative context, and use them to prompt LVLMs to answer user queries.
+For comparison, we obtain standard responses via regular decoding.
+We then use QwenVL-Max-0125 Bai et al.
+(2023b) to extract image-related words from the generated responses.
+Inspired by CoT Wei et al.
+(2022), we guide the model through a multi-stage extraction process: it first identifies all visually relevant words without access to the image, and then selects those strongly associated with the image content, which we define as relevant words.
+Finally, we compute the overlap ratio between the relevant words generated from negative contexts and those obtained through regular decoding.
+This ratio quantifies how much essential visual content is preserved in responses generated from negative contexts, serving as an indicator of retained visual information.
+The prompt used to guide QwenVL-Max is shown in Fig.
+9.
+Based on this analysis, Fig.
+2 shows that VCD and ICD retain over 30% of essential visual information, which is substantially weakened during contrastive decoding.
+In contrast, our method results in a negligible amount of residual information, effectively preventing visual information loss.
+Figure 8: The results on the validation set of CHAIR using LLaVA-1.5. The X-axis represents the value of $\gamma$.
+Appendix E Case Study
+We present case studies to illustrate the superior performance of our method in hallucination reduction, as shown in Fig.
+10 and Fig.
+11.
+The image captions are generated by LLaVA-1.5.
+Red highlights indicate hallucinated content, while green highlights denote content consistent with the image.
+We can clearly observe that Regular Decoding produces a substantial amount of hallucinated content, often manifesting as content words such as objects, attributes, and relationships.
+In Fig.
+10, for example, although the image depicts a skateboarding scene, the caption includes a reference to "spectators"—a detail not present in the image.
+Similarly, Fig.
+11 shows an eating scene, yet the caption introduces a "long dining table," which is not actually visible.
+These hallucinations stem from LVLMs inheriting prior knowledge from LLMs, leading to biased descriptions in certain typical scenarios.
+In contrast, our method significantly reduces such hallucinations.
+While Fig.
+11 still contains a hallucinated object (chips), this instance can be attributed to image ambiguity, where even human observers might plausibly infer such content.
+Overall, our method effectively mitigates the influence of language priors by leveraging cross-image consistency.
+Figure 9: Prompt for guiding QwenVL-Max to progressively extract relevant words from the response.
+Figure 10: Case study (1).
+Figure 11: Case study (2).

### v5

@@
 Abstract
-Language priors are a major cause of hallucinations in Large Vision-Language Models (LVLMs), often leading to text that is linguistically plausible but visually inconsistent.
-Recent work explores contrastive decoding as a training-free solution, but these methods typically construct negative contexts from the original image, resulting in visual information loss and distorted distribution.
-Motivated by the observation that language priors stem from the LLM backbone and remain consistent across images, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective training-free method that uses different images to construct negative contexts.
-We further analyze the cross-image behavior of language priors and introduce a distinction between essential priors (supporting fluency) and detrimental priors (causing hallucinations).
-By selectively preserving essential priors and suppressing detrimental ones, our method reduces hallucinations while maintaining coherent and fluent language generation.
-Experiments on 4 benchmarks and 6 LVLMs across three model families confirm the effectiveness and generalizability of CICD, especially in image captioning, where language priors are particularly pronounced.
-Code will be released once accepted.
+Over-reliance on language priors is a major cause of hallucinations in Large Vision-Language Models (LVLMs), often leading to outputs that are linguistically plausible but visually inconsistent.
+Recent studies have explored contrastive decoding as a training-free solution.
+However, these methods typically construct contrastive visual inputs by perturbing the original image, resulting in distorted contrastive distributions, incomplete contrastive signals, and excessive suppression of language priors.
+Motivated by the observation that language priors tend to remain consistent across different images, we propose Cross-Image Contrastive Decoding (CICD), a simple yet effective training-free method that uses unrelated images as contrastive visual inputs.
+To address the issue of over-suppressing language priors, which can negatively affect the quality of generated responses, we further introduce a dynamic selection mechanism based on the cross-image differences in model behavior.
+By selectively suppressing language priors, our method reduces hallucinations without compromising the model’s performance.
+Extensive experiments across multiple benchmarks and LVLMs confirm the effectiveness and generalizability of CICD, particularly in image captioning, where language priors are especially dominant.
+[^1] Our code is available at https://github.com/beta-nlp/CICD.
 Cross-Image Contrastive Decoding: Precise, Lossless Suppression of Language Priors in Large Vision-Language Models
-Jianfei Zhao1,2, Feng Zhang1, Xin Sun1, Chong Feng1,3,thanks: Corresponding Autor: fengchong@bit.edu.cn , 1School of Computer Science and Technology, Beijing Institute of Technology, Beijing, China, 2Zhongguancun Academy, Beijing, China, 3Southeast Academy of Information Technology, Beijing Institute of Technology, Fujian, China,
+Jianfei Zhao1,2, Feng Zhang1, Xin Sun1, Lingxing Kong4, Zhixing Tan4, Chong Feng1,3 1School of Computer Science and Technology, Beijing Institute of Technology 2Zhongguancun Academy 3Southeast Academy of Information Technology, Beijing Institute of Technology 4Tsinghua University {zhqingan, bit_zhangfeng, sunxin}@bit.edu.cn, klxstar@126.com, tzx.2019@tsinghua.org.cn, fengchong@bit.edu.cn
 1 Introduction
-Figure 1: A case of language prior. Adding the same prefix when captioning another image will lead to the same language priors. The model used is LLaVA-1.5.
-Recent advances in Large Language Models (LLMs) Brown et al.
-(2020); Touvron et al.
-(2023); Chiang et al.
-(2023); Bai et al.
-(2023a) have opened promising pathways toward Artificial General Intelligence (AGI).
-Building on these developments, numerous studies Liu et al.
+Recent developments in Large Language Models (LLMs) Chiang et al.
+(2023a); Bai et al.
+(2023a) have opened a promising pathway toward Artificial General Intelligence (AGI).
+Building on these advances, many studies Liu et al.
 (2023); Dai et al.
 (2023); Bai et al.
-(2023b, 2025) have extended LLMs by integrating visual inputs, giving rise to Large Vision-Language Models (LVLMs) capable of understanding visual content.
-They are capable of addressing a broad spectrum of vision-language tasks, such as image captioning Jing et al.
-(2024); Rohrbach et al.
+(2023b) have focused on integrating visual inputs into LLMs, giving rise to Large Vision-Language Models (LVLMs) capable of understanding visual content.
+These models can handle a wide range of vision-language tasks, including image captioning Rohrbach et al.
 (2018); Ye et al.
 (2025); Chen et al.
-(2025), visual question answering (VQA) Lovenia et al.
-(2024); Li et al.
-(2023b); Yu et al.
-(2024), and visual grounding Rasheed et al.
-(2024); You et al.
-(2024); Lu et al.
+(2025) and visual question answering (VQA) Li et al.
+(2023b); Lovenia et al.
+(2024); Yu et al.
 (2024).
-Despite impressive capabilities, LVLMs remain susceptible to hallucinations Li et al.
-(2023c); Wang et al.
-(2023), often generating text that contradicts the visual input.
-These hallucinations stem from various sources, with language priors Chen et al.
-(2025); Leng et al.
+Figure 1: Comparison of contrastive decoding–based methods.
+Our approach avoids overlapping visual information, training-inference gap, and excessive suppression of language priors observed in existing methods.
+Despite their impressive capabilities, LVLMs are known to suffer from severe hallucinations Li et al.
+(2023b); Wang et al.
+(2023b); Leng et al.
+(2024).
+These models often generate responses based solely on memorized linguistic associations rather than actual visual input.
+For example, when describing an office scene, an LVLM may erroneously mention a “computer” even if no such object is present in the image.
+Therefore, improving the reliability and applicability of LVLMs across domains requires addressing this hallucination problem.
+While the causes of hallucination in LVLMs are complex, prior studies Leng et al.
 (2024); Wang et al.
-(2024c); Zhu et al.
-(2024b) being the most fundamental underlying cause.
-Specifically, since LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge encoded in the LLMs.
-As a result, their outputs are influenced not only by visual evidence but also by pre-learned language patterns.
-This often leads to a tendency to rely on memorized linguistic associations when responding to user queries.
-For instance, when describing an office scene, an LVLM may erroneously mention a "computer" even if no such object is present in the image.
-Figure 2: The overlap ratio of image-related words between responses generated from negative contexts and those from regular inputs.
-Both VCD and ICD exhibit over 30% overlap, indicating that their negative contexts retain a considerable amount of essential visual information.
-In contrast, our method results in a nearly negligible overlap.
-The detailed analysis methodology can be found in Appendix.
-D.
-Methods for mitigating language priors in LVLMs can generally be classified into two categories.
-The first involves constructing specialized training datasets to reduce the model’s dependence on linguistic biases Zhibo et al.
-(2023); Wu et al.
+(2024b); Zhu et al.
+(2025a); Chen et al.
+(2025) have identified over-reliance on language priors as one of the primary contributing factors.
+Recent efforts to mitigate this issue have focused on plug-and-play, training-free methods Wang et al.
+(2024b); Leng et al.
+(2024); Zhu et al.
+(2025a); Zhang et al.
+(2025), which avoid the substantial computational costs associated with training-based approaches Wu et al.
+(2022); Ren et al.
 (2023); Chen et al.
-(2025), though such approaches often incur substantial training costs.
-Another line of work explores plug-and-play, training-free methods, which are also the focus of our study.
-Among them, many recent studies adopt contrastive decoding Li et al.
-(2023a), which constructs negative contexts to reveal and suppress language priors.
-Some approaches directly perturb the input image Leng et al.
+(2025).
+These methods Leng et al.
 (2024); Wang et al.
-(2024c) to alter the visual context, while others enhance the visual features extracted from the original image Zhu et al.
+(2024b); Zhu et al.
+(2025a) commonly employ contrastive decoding Li et al.
+(2023a), which highlights differences in output distributions between the original and contrastive inputs.
+Various strategies Leng et al.
+(2024); Wang et al.
+(2024b); Liu et al.
+(2024d); Favero et al.
+(2024); Zhang et al.
+(2025) have been proposed for constructing visual inputs in contrastive decoding.
+While previous studies have demonstrated their effectiveness, we identify three major challenges in contrastive decoding-based methods, as illustrated in Figure 1.
+First, distorted distributions from contrastive visual inputs.
+Leng et al.
+(2024); Wang et al.
+(2024b); Liu et al.
+(2024d); Favero et al.
+(2024) propose using noised images or even removing the image entirely when computing the contrastive distribution.
+However, these approaches introduce a training-inference gap, as such contrastive inputs are not seen during training.
+This may result in distorted contrastive distributions.
+Second, overlapping visual information between original and contrastive inputs.
+Leng et al.
+(2024); Wang et al.
 (2024b); Zhang et al.
+(2025) propose using noisy or generated images that share visual similarities with the original inputs.
+However, such strategies may lead to incomplete contrastive signals.
+Third, trade-offs in suppressing language priors.
+Although over-reliance on language priors can cause hallucinations, excessively suppressing them may degrade the performance of the model.
+Although Zhang et al.
+(2025) selectively remove language priors, they augment the retained language priors, which introduces bias.
+As a result, mitigating language priors through contrastive decoding remains an open challenge that needs further investigation.
+Figure 2: Adding the same prefix when captioning a different image leads to the same hallucination, indicating that language priors are consistent across images.
+In this paper, we propose Cross-Image Contrastive Decoding (CICD), a simple yet effective method to address the aforementioned challenges in contrastive decoding.
+We begin by observing that LVLMs often produce highly similar outputs when given the same textual prefix but entirely different images, suggesting that language priors tend to dominate regardless of visual input (see Figure 2 for an example).
+Based on this observation, we construct contrastive visual inputs using different images.
+This approach avoids the training-inference gap and minimizes overlap between the original and contrastive inputs, thereby emphasizing the visual information in the original input and alleviating over-reliance on language priors.
+To prevent excessive suppression of language priors during contrastive decoding, we leverage the consistency of model behavior across images to dynamically determine, at each decoding step, whether language priors should be suppressed and to what extent.
+When the divergence falls below a predefined threshold, our method backs off to regular decoding.
+This adaptive mechanism allows CICD to effectively reduce over-reliance on language priors while preserving the model’s performance.
+We conduct extensive experiments across multiple benchmarks and LVLMs. The results demonstrate that our method effectively alleviates language priors, particularly in the image captioning task.
+2 Related Work
+2.1 Hallucination in LVLMs
+In the multimodal domain, the study of hallucinations primarily focuses on cross-modal consistency Li et al.
+(2023b); Wang et al.
+(2023b), which requires that content generated by LVLMs remains faithful to the visual input.
+Multiple factors contribute to hallucinations in LVLMs, including biased visual attention Zhu et al.
+(2025b); Li et al.
+(2025), attention misallocation Huang et al.
+(2024); Yin et al.
+(2025); Kang et al.
+(2025), and over-reliance on language priors Leng et al.
+(2024); Zhu et al.
+(2025a); Wang et al.
+(2024b).
+Our work primarily focuses on mitigating the over-reliance on language priors in LVLMs.
+2.2 Language Priors in LVLMs
+As LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge embedded in LLMs—commonly referred to as language priors Wu et al.
+(2022); Ren et al.
 (2023).
-These methods typically compute the difference between the logit distributions obtained from the original and perturbed (i.e., negative) contexts to reduce the model’s reliance on language priors.
-Overall, following the paradigm introduced in VCD Leng et al.
-(2024), these methods routinely perturb the original inputs to generate negative contexts.
-However, this prevailing design choice raises a critical question: Is perturbing the original context to construct negative contexts truly the most effective strategy for inducing and mitigating language priors?
-Figure 3: Essential priors (left) and detrimental priors (right).
-We randomly select 500 images from the MSCOCO dataset and visualize LLaVA-1.5’s output by the t-SNE algorithm van der Maaten and Hinton (2008).
-w/ Image represents regular outputs.
-w/o Image represents the outputs after removing the input image.
-w/ Neg.
-Image represents replacing the input image by the most dissimilar image in the dataset retrieved by CLIP Radford et al.
+Over-reliance on these priors can cause LVLMs to generate content that is linguistically plausible but inconsistent with visual evidence.
+To mitigate this issue, Liu et al.
+(2024a); Gunjal et al.
+(2024); Chen et al.
+(2025) design specialized training datasets.
+However, these training-based approaches are often resource-intensive and not easily transferable to broader application scenarios.
+Recent studies have explored contrastive decoding Li et al.
+(2023a); Chuang et al.
+(2023) as a training-free alternative to reduce dependence on language priors.
+These methods construct an alternative logit distribution alongside the original one, and then contrast the two distributions to suppress language priors during decoding.
+The contrastive distribution is typically generated using techniques such as masking the image Leng et al.
+(2024), perturbing the instruction Wang et al.
+(2024b), augmenting the visual input Zhu et al.
+(2025a), or performing cross-modal conversion Zhang et al.
 (2025).
-We argue that the answer is unequivocally negative.
-Constructing negative contexts based on the original inputs can lead to two main issues: (1) Visual information loss, where the negative context retains partial visual information, which may be degraded during contrastive decoding.
-As shown in Fig.
-2, compared to regular responses, those generated from negative contexts retain over 30% of the essential visual information, which is subsequently weakened during the contrastive decoding process.
-(2) Distorted distribution, where perturbing the model’s input to induce priors may disrupt its internal representations, ultimately resulting in a distorted logit distribution.
-This analysis raises an important question: How should negative contexts be constructed to avoid the above issue?
-Figure 4: A case of JS divergence between the logit distributions of two images.
-In fact, simply taking a different image to construct negative context is sufficient.
-This insight is based on the thought that the language priors guiding step-wise token generation in LVLMs originate from the LLM backbone and are largely independent of the input image.
-As shown in Fig.1, LVLM produces highly consistent outputs when given completely different images but the same prefix, which strongly supports our hypothesis.
-Building on this, we propose Cross-Images Contrastive Decoding (CICD), a simple yet effective method for suppressing language priors in LVLMs.
-Instead of perturbing the original inputs, CICD constructs the negative context using a different image that shares the same language priors.
-Our method avoids the two key issues present in previous approaches.
-Since the overlap between the original and the new image is limited, our approach avoids introducing critical visual information into the negative context.
-As shown in Fig.
-2, only a negligible amount (approximately 0.11%) of essential visual information is retained in our negative context.
-Moreover, it enables the model to reveal intrinsic priors, maintaining the model’s natural inference process without injecting external bias.
-The formal analysis of our method is presented in Sec.
-3.2.
-Focusing on language priors, we further explore whether all such priors should be strictly eliminated.
-The answer is intuitively negative.
-Language priors can be divided into two categories: (1) Essential Priors, which are necessary for maintaining textual fluency, and (2) Detrimental Priors, which introduce information unrelated to the visual input.
-Clearly, only the detrimental priors should be removed.
-This raises a critical question: How can we accurately distinguish between the two and selectively eliminate only the detrimental ones?
-As shown in Fig.3, essential priors exhibit high consistency across different images, whereas detrimental priors show substantial divergence when generating image-related content.
-Based on this observation, we propose to distinguish essential from detrimental priors using cross-image consistency, as illustrated in Fig.4.
-Specifically, we compute the Jensen–Shannon (JS) divergence between the logit distributions of different visual contexts, and apply a stable threshold to identify essential priors that should be retained.
-Extensive experiments on 4 widely used benchmarks and 6 LVLMs across three model families demonstrate the effectiveness and generalizability of our method.
-By suppressing language priors, CICD consistently reduces hallucinations and achieves state-of-the-art performance across all benchmarks.
-The improvement is particularly pronounced in the image captioning task, which tends to suffer more from language priors due to its longer textual outputs.
-In summary, our main contributions are as follows:
-  • We analyze existing contrastive decoding methods in depth and reveal their limitations stemming from reliance on the original image.
-To address this, we verify the consistency of language priors across images and propose a stable approach to distinguish essential priors from detrimental ones.
-  • Based on this analysis, we propose CICD (Cross-Images Contrastive Decoding), which leverages retrieved images to induce language priors while selectively preserving essential priors during the mitigation process.
-  • Extensive experiments on multiple benchmarks and LVLMs demonstrate the superiority of our method in alleviating language priors, particularly in the image captioning task.
-2 Related Work
-Due to space limitations, we provide a brief overview here. A more detailed discussion of related work can be found in Appendix.A.
-2.1 Hallucination in LVLMs
-Hallucinations in LVLMs primarily concern cross-modal consistency Li et al.
-(2023c); Wang et al.
-(2023), which requires that the generated content remains faithful to the visual input.
-Among the various causes of hallucination, language priors Zhibo et al.
-(2023); Wu et al.
-(2022) are the most fundamental.
-These priors originate from the LLM’s pretraining corpus and guide generation based on learned language patterns, which can produce fluent yet visually inconsistent outputs.
-Several approaches Chen et al.
-(2025); Liu et al.
-(2024a); Gunjal et al.
-(2024) attempt to mitigate language priors by constructing specialized training datasets, but these methods often incur significant additional training costs.
-Alternatively, recent training-free methods have explored contrastive decoding Li et al.
-(2023a); Chuang et al.
-(2024), typically constructing negative contexts based on the original image Leng et al.
-(2024); Wang et al.
-(2024c); Zhu et al.
-(2024b); Zhang et al.
-(2025) in an attempt to elicit and suppress language priors.
-In this work, we identify critical limitations of these approaches that stem from their reliance on modifying the original image.
-Specifically, such modifications may lead to visual information loss and distorted distribution.
-We further conduct an in-depth analysis of the cross-image consistency of language priors and, based on this insight, propose CICD.
-Additionally, we introduce a method for distinguishing between essential priors and detrimental priors, enabling fine-grained and precise mitigation of language priors in LVLMs.
-3 Preliminary
-Figure 5: The diagram of our method.
+Our approach also builds on the contrastive decoding framework.
+However, unlike the aforementioned methods, we propose constructing contrastive inputs using a different image and introducing a dynamic mechanism to prevent the over-suppression of language priors.
+This design avoids both the introduction of overlapping visual information and the training-inference discrepancies in the contrastive input, making it more effective at addressing hallucinations in LVLMs.
+3 Cross-Image Contrastive Decoding
+We propose Cross-Image Contrastive Decoding (CICD) for effective mitigation of language priors.
+CICD first retrieves a distinct image as the contrastive visual input.
+Then, at each decoding step, it dynamically chooses between regular decoding and contrastive decoding based on the divergence between the original and contrastive distributions.
+An overview of CICD is presented in Figure 3.
+We begin with an analysis of language priors in Section 3.1.
+Based on this analysis, we present methods for constructing contrastive visual inputs in Section 3.2.
+Finally, we discuss the importance of balanced language prior suppression in Section 3.3.
+Figure 3: Overview of our method.
 We input the target image and a distinct image into the LVLM separately.
-During the step-by-step generation of the response, we compute the Jensen-Shannon (JS) divergence between the two logit distributions at each step to distinguish essential priors from detrimental ones.
-For detrimental priors, we apply contrastive decoding to suppress harmful language biases, while regular decoding is used for essential priors.
-3.1 Contrastive Decoding
-LVLMs take as input a concatenation of vision tokens from a vision encoder and an encoded text prompt.
-After forward propagation, the model projects the output into a logit distribution to compute conditional probabilities and generate responses in an autoregressive manner:
-$y_{t}\sim p_{\theta}(y_{t}\mid v,x,y_{<t})$ (1)
-$=$ $\ \mathrm{softmax}\left[\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})\right]$
-Here, $v$ and $x$ represent the visual and textual inputs, respectively. $\mathrm{logit}_{\theta}$ represents the logit distribution formulated by the model with parameters $\theta$.
-Since LVLMs adopt pretrained LLMs as their backbone, they inherently inherit language priors independent of the visual input.
-As a result, tokens ranked higher in the output distribution $\mathrm{logit}_{\theta}$ may appear semantically plausible but can be objectively inconsistent with the visual evidence.
-This constitutes a fundamental cause of hallucinations.
-A commonly used approach to mitigate such priors is contrastive decoding Li et al.
-(2023a); Leng et al.
-(2024), which modifies the original image to construct a native visual context that elicits language priors.
-These priors are then suppressed by subtracting the logit distribution derived from the native context from that of the original context.
-$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,x,y_{<t})$ (2)
-$=$ $(1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})-\alpha\,\mathrm{logit}_{N}$
-Here, $\mathrm{logit}_{N}$ denotes the logit distribution obtained from the native context, and $\mathrm{logit}_{\mathrm{CD}}$ represents the final distribution used for next-token prediction.
-3.2 Formal Analysis
-The core component of contrastive decoding lies in the construction of the native context.
-Our preliminary investigation reveals that existing methods, which modify the original image, introduce several limitations (see Sec.1).
-Further observations and analysis show that language priors are inherent characteristics of LVLMs and remain consistent across different images, as illustrated in Fig.1.
-This insight leads to a natural alternative, where the native context is constructed using an image distinct from the original.
-Under this design, Eq.
-2 is refined as:
-$\mathrm{logit}_{\mathrm{CD}}(y_{t}\mid v,x,y_{<t})$ (3)
-$=$ $\ (1+\alpha)\,\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$
-$\quad-\alpha\,\mathrm{logit}_{\theta}(y_{t}\mid v^{\prime},x,y_{<t})$
-where $v\textquoteright$ ’ is a different image from $v$. We further provide a formal analysis of the effectiveness of this approach.
-Most existing methods employ sampling decoding, where a token is randomly selected according to the logit distribution.
-We define the set of plausible tokens under a given logit distribution as $S=V\cup T$, where $V$ represents tokens related to visual content, and $T$ represents tokens unrelated to the image, i.e., those driven by language priors.
-The negative context constructed by a different image yields a new logit distribution, whose plausible token set is $S\textquoteright=V\textquoteright\cup T$, given that language priors remain unchanged across images.
-The contrastive decoding operation can thus be interpreted as subtracting one plausible token set from another:
-$S-S^{\prime}=(V\cup T)-(V^{\prime}\cup T)=V\mbox{ if }V^{\prime}\cap V=\emptyset$ (4)
-In other words, this subtraction removes all language priors that are unrelated to the visual input.
-This approach achieves three objectives simultaneously: (1) retention of all essential visual information, (2) precise elimination of language priors, and (3) compatibility with the input distribution expected by LVLMs.
-4 Method
-Leveraging the cross-image consistency of language priors, we propose CICD (Cross-Images Contrastive Decoding), which effectively mitigates language priors while preventing visual information loss.
-In addition, we retain essential linguistic priors to ensure fluent and natural generation.
-As illustrated in Fig.
-5, CICD comprises three key steps: (1) retrieving a distinct image, (2) identifying detrimental language priors, and (3) removing these priors through contrastive decoding.
-4.1 Image Retrieval
-As discussed earlier, it is essential to select an image with limited information that overlaps with the input image.
+During step-by-step response generation, we compute the Jensen–Shannon (JS) divergence between the two logit distributions at each step to decide whether to suppress language priors.
+Contrastive decoding is applied to suppress language priors, while regular decoding is used to avoid excessive suppression.
+3.1 An Analysis of Language Priors
+Figure 4: Similarity of the output logit distributions of LLaVA-1.5-7B in image-relevant and image-irrelevant contexts across 500 images.
+While over-reliance on language priors may lead to hallucinations in LVLMs, excessive suppression of these priors can negatively affect the quality of generated texts.
+To investigate the effect of language priors on the generation of image-relevant and image-irrelevant tokens, we examine the logit distributions of LVLMs in both image-relevant and image-irrelevant contexts.
+To this end, we randomly selected 500 images from the MSCOCO dataset and compared the similarity of the output logit distributions produced by LLaVA-1.5-7B in image-relevant and image-irrelevant contexts.
+We define the generation for the first token as the image-irrelevant context, and use the prefix “The picture depicts a” as the image-relevant context.
+We use cosine distance, Euclidean distance, and variation to study the similarity between logit distributions in the image-irrelevant and image-relevant contexts, respectively.
+It is evident that the logit distributions in image-irrelevant contexts exhibit high similarity across different images, regardless of the similarity metric used.
+In contrast, the logit distributions in image-relevant contexts do not display such consistency.
+We further validate this observation using additional context prompts, and the results remain consistent, suggesting that cross-image similarity can be leveraged to discriminate between image-irrelevant and image-relevant contexts.
+Only language priors in image-relevant contexts need to be removed.
+3.2 Construction of Contrastive Visual Inputs
+Based on the above analysis, we propose constructing contrastive visual inputs using an image that is distinct from the original input image.
+By choosing an image from a given dataset instead of synthesizing one, we avoid introducing a potential training–inference gap while maximizing visual contrast during contrastive decoding.
+Intuitively, the image with the lowest similarity to the input image is preferred as the contrastive visual input for contrastive decoding.
+We propose two different methods for selecting a distinct image from a given dataset $\mathcal{D}_{\textrm{img}}$:
+Retrieval.
 Since visual features are extracted by the vision encoder, a natural strategy is to measure the similarity between two images using the encoder itself to estimate their overlap.
-In other words, the image with the lowest similarity to the input image is selected as the reference.
 To ensure better generalizability, we adopt the pretrained CLIP model Radford et al.
 (2021) as the retriever.
-All images in the dataset are vectorized in advance, and during inference, the image with the lowest cosine similarity to the input is retrieved as the contrastive image.
-Although vector retrieval is relatively efficient given the computational cost of LVLMs, practical deployment still requires loading an additional retriever and maintaining a vector database.
-As illustrated in Fig.1, we observe that language priors primarily manifest in content words such as objects, attributes, and relationships.
+As all images in the dataset $\mathcal{D}_{\mathrm{img}}$ can be vectorized in advance before the inference stage, the retrieval process incurs negligible computational cost.
+We simply retrieve the image with the lowest cosine similarity to the input, which is then used as the contrastive image during decoding.
+Random.
+Although vector retrieval is relatively efficient compared to the computational cost of LVLMs, practical deployment still requires loading an additional retriever and maintaining a vector database.
+Previous works Li et al.
+(2023b); Wang et al.
+(2023a); Ye et al.
+(2025) reveal that language priors primarily exist in content words such as objects, attributes, and relationships.
 These visual elements tend to exhibit minimal overlap across natural images.
 Therefore, randomly selecting an image can serve as an effective and lightweight alternative for practical applications.
-We provide further discussion in Appendix.C.2.
-Note that, to ensure reproducibility, we adopt the retrieval-based strategy for consistent and deterministic image selection throughout all experiments.
-4.2 Identifying Detrimental Priors
-Although language priors may cause LVLMs to generate hallucinated content, they remain essential for producing fluent and coherent text.
-To address this, we categorize language priors into two types: essential priors and detrimental priors, with only the latter requiring elimination.
-Essential priors are defined as purely linguistic patterns that are entirely independent of visual input.
-As discussed in Sec.
-1, the logit distributions associated with essential priors should remain highly consistent across diverse visual inputs.
-This intrinsic property provides a reliable basis for distinguishing between essential and detrimental priors.
+3.3 Dynamic Language Prior Suppression
+Although language priors may cause LVLMs to generate hallucinated content, they are essential for the model’s generation capabilities.
+As previously analyzed, language priors exhibit high consistency across different images in image-irrelevant contexts.
+Therefore, we employ the Jensen–Shannon (JS) divergence to measure the consistency between two logit distributions produced from different images $\bm{v}$ and $\bm{v}^{\prime}$, which is formally defined as
+$d(\bm{v},\bm{v}^{\prime})=\mathrm{JSD}(p_{\bm{\theta}}(y_{t}|\bm{v},\mathbf{x},\mathbf{y}_{<t})\ \|\ p_{\bm{\theta}}(y_{t}|\bm{v}^{\prime},\mathbf{x},\mathbf{y}_{<t}),$ (1)
+where $\mathrm{JSD}(P\|Q)=\frac{1}{2}\mathrm{KL}(P\|M)+\frac{1}{2}\mathrm{KL}(Q\|M)$, $M=\frac{1}{2}(P+Q)$, $\mathrm{KL}(P\|Q)$ denotes the KL divergence between two probability distributions $P$ and $Q$, and $p_{\bm{\theta}}(y_{t}\mid\bm{v},\mathbf{x},\mathbf{y}_{<t})$ is the probability distribution of the next token $y_{t}$ given visual input $\bm{v}$, textual input $\mathbf{x}$, and the previously generated token sequence $\mathbf{y}_{<t}$.
+A higher value of $d(\bm{v},\bm{v}^{\prime})$ indicates that the current context is more image-relevant, while a lower value may suggest an image-irrelevant generation.
+Figure 5: JS divergence between the logit distributions of two distinct images.
 Figure 6: 500 random images from the MSCOCO dataset captioned by LLaVA-1.5.
 For each next-token prediction, we statistic the JS divergence (JSD) between two logit distributions produced by different images.
-We employ the Jensen–Shannon (JS) divergence to measure the consistency between two logit distributions produced from different images:
-$\mathcal{D}_{\mathrm{JS}}=\mathrm{JS}\left(p_{\theta}(y_{t}\mid v,x,y_{<t})||p_{\theta}(y_{t}\mid v\textquoteright,x,y_{<t})\right)$ (5)
-where $p_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq.1.
-The statistical results are visualized in Fig.6.
-We observe that the JSD values can be clearly separated into two distinct ranges: $\log_{10}(\mathcal{D}_{\mathrm{JS}})>-4$ and $\log_{10}(\mathcal{D}_{\mathrm{JS}})\leq-4$.
-Higher JSD values are associated with detrimental priors, while lower values correspond to essential priors, as shown in Fig.3 and Fig.4.
-Based on this observation, we set a threshold of $\gamma=-4$ to distinguish between the two types of priors.
-Further discussion can be found in Appendix.
-C.3
+To demonstrate the effectiveness of JS divergence in determining context relevance, we plot the logarithm of the JS divergence between distributions generated with different images at each decoding step.
+Figure 5 shows the results.
+Clearly, image-relevant tokens such as “_people”, “_white”, and “_red” exhibit higher divergence values (often above 4-4 when taking $\log_{10}(d(\bm{v},\bm{v}^{\prime}))$), while function words such as “_of”, “_be” and “_the” have lower JS divergence values (often below 4-4 in $\log_{10}(d(\bm{v},\bm{v}^{\prime}))$).
+We further randomly select 500 images from the MSCOCO dataset, captioned by LLaVA-1.5, and compute the JS divergence between logit distributions produced by different images for each next-token prediction.
+The statistical results are visualized in Figure 6.
+We observe that the JS divergence values also fall into two distinct ranges: $\log_{10}(d(\bm{v},\bm{v}^{\prime}))>-4$ and $\log_{10}(d(\bm{v},\bm{v}^{\prime}))\leq-4$.
+These results further confirm that JS divergence can effectively distinguish between image-relevant and image-irrelevant contexts.
+Based on the above analysis, CICD enables dynamic suppression of language priors.
+Specifically, CICD performs regular decoding at steps where image-irrelevant contexts are detected, while applying contrastive decoding with different images to suppress language priors in image-relevant contexts.
+A hyperparameter $\gamma$ is used to determine whether the current decoding step is image-irrelevant.
+Formally, the final logit in CICD is calculated as
+$\mathrm{logit}(y_{t})$ $=\left\{\begin{array}[]{ll}\mathrm{logit}_{\theta}(y_{t}\ |\ \bm{v},\mathbf{x},\mathbf{y}_{<t}),&\textrm{if }d(\bm{v},\bm{v}^{\prime})\leq\gamma\\ \mathrm{logit}_{\textrm{CD}}(y_{t}\ |\ \bm{v},\mathbf{x},\mathbf{y}_{<t}),&\textrm{if }d(\bm{v},\bm{v}^{\prime})>\gamma\end{array}\right$.
+(4)
+where $\mathrm{logit}_{\textrm{CD}}(y_{t}|\bm{v},\mathbf{x},\mathbf{y}_{<t})$ is defined as
+$\mathrm{logit}_{\textrm{CD}}(y_{t}|\bm{v},\mathbf{x},\mathbf{y}_{<t})=$ $(1+\alpha)\textrm{logit}_{\bm{\theta}}(y_{t}|\bm{v},\mathbf{x},\mathbf{y}_{<t})$ (5)
+$-\alpha\textrm{logit}_{\bm{\theta}}(y_{t}|\bm{v}^{\prime},\mathbf{x},\mathbf{y}_{<t}),$
+where $\alpha$ is the contrastive coefficient.
+Unlike previous approaches such as Leng et al.
+(2024), $\alpha$ is not treated as a hyperparameter in our method.
+Instead, it is dynamically set to $1-\log_{10}(d(\bm{v},\bm{v}^{\prime}))$.
+This is based on the intuition that detrimental priors with higher similarity are more difficult to remove and therefore require stronger suppression.
+To ensure numerical stability, we clip $\alpha$ to the interval $[1,3]$.
 Setting Method LLaVA-1.5 InstructBLIP Qwen-VL-Chat
-Acc. $\uparrow$ F1$\uparrow$ Acc. $\uparrow$ F1$\uparrow$ Acc. $\uparrow$ F1$\uparrow$
+[…]
 Random Regular 83.47 82.25 80.49 81.10 82.32 79.35
-VCD 85.30 84.59 81.75 82.06 84.74 82.80
-ICD 84.81 83.44 80.95 81.59 83.10 80.52
-IBD 85.36 84.28 74.82 77.87 83.07 80.23
-DeFG 86.68 85.71 84.35 84.50 84.04 81.49
-CICD 87.85 87.77 84.00 84.89 87.54 86.37
+VCD Leng et al. (2024) 85.30 84.59 81.75 82.06 84.74 82.80
+ICD Wang et al. (2024b) 84.81 83.44 80.95 81.59 83.10 80.52
+PAI Liu et al. (2024d) 84.56 83.33 79.54 80.27 84.14 81.77
+IBD Zhu et al. (2025a) 85.36 84.28 74.82 77.87 83.07 80.23
+DeGF Zhang et al. (2025) 86.68 85.71 84.35 84.50 84.04 81.49
+CICD (ours) 87.85 87.77 84.00 84.89 87.54 86.37
 Popular Regular 79.56 78.99 76.59 78.16 81.11 78.34
-VCD 80.81 80.84 77.22 78.60 82.95 81.17
-ICD 81.44 80.53 76.82 78.48 81.26 78.87
-IBD 81.68 81.13 69.35 74.31 82.20 79.59
-DeFG 82.96 82.46 79.25 80.46 83.02 80.57
-CICD 84.83 85.18 81.01 82.53 85.94 84.90
+VCD Leng et al. (2024) 80.81 80.84 77.22 78.60 82.95 81.17
+ICD Wang et al. (2024b) 81.44 80.53 76.82 78.48 81.26 78.87
+PAI Liu et al. (2024d) 80.27 79.72 75.04 77.05 83.06 80.89
+IBD Zhu et al. (2025a) 81.68 81.13 69.35 74.31 82.20 79.59
+DeGF Zhang et al. (2025) 82.96 82.46 79.25 80.46 83.02 80.57
+CICD (ours) 84.83 85.18 81.01 82.53 85.94 84.90
 Adversarial Regular 76.95 77.13 72.89 75.58 78.23 75.72
-VCD 77.27 78.19 74.03 76.39 80.37 78.96
-ICD 78.07 77.98 73.19 75.91 79.03 77.05
-IBD 78.26 78.52 66.60 72.63 79.68 77.36
-DeFG 79.40 79.69 75.19 77.52 80.42 78.24
-CICD 78.60 80.30 75.55 78.63 81.94 81.37
+VCD Leng et al. (2024) 77.27 78.19 74.03 76.39 80.37 78.96
+ICD Wang et al. (2024b) 78.07 77.98 73.19 75.91 79.03 77.05
+PAI Liu et al. (2024d) 77.02 77.30 72.14 75.15 80.62 78.76
+IBD Zhu et al. (2025a) 78.26 78.52 66.60 72.63 79.68 77.36
+DeGF Zhang et al. (2025) 79.40 79.69 75.19 77.52 80.42 78.24
+CICD (ours) 78.60 80.30 75.55 78.63 81.94 81.37
 Table 1: Results on POPE.
@@
 The bolded results denote the best in each setting, and the underlined results denote the second-best.
-4.3 Cross-Images Contrastive Decoding
-After completing the above preparation, we integrate the proposed method into the contrastive decoding process.
-Specifically, the model performs regular decoding when essential priors are detected and applies contrastive decoding to suppress detrimental priors.
-In the case of detrimental priors, smaller distributional differences imply a stronger detrimental effect.
-Accordingly, we define the contrastive coefficient $\alpha$ in Eq.
-3 as: $\alpha=1-\log_{10}(\mathcal{D}_{\mathrm{JS}})$.
-Based on the observed density of $\mathcal{D}_{\mathrm{JS}}$ values, we clip $\alpha$ to the interval $[1,3]$ to ensure numerical stability.
-The final logit distribution is defined as follows:
-$\mathrm{logit}_{CICD}(y_{t}\mid v,x,y_{<t})=$ (6)
-$\begin{cases}\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}\leq\gamma\\ \mathrm{logit}_{CD}(y_{t}\mid v,x,y_{<t})&\text{if }\mathcal{D}_{\mathrm{JS}}>\gamma\end{cases}$
-where $\mathrm{logit}_{\theta}(y_{t}\mid v,x,y_{<t})$ is defined in Eq. 1, and $\mathrm{logit}_{CD}(y_{t}\mid v,x,y_{<t})$ is defined in Eq. 3.
 4 Experiment
+4.1 Setup
+Benchmarks.
+We evaluate our CICD on 4 widely adopted multimodal hallucination benchmarks: DetailCaps Ye et al.
+(2025): A fine-grained image captioning benchmark.
+We evaluate on a 700-image MSCOCO subset.
+POPE Li et al.
+(2023b): A hallucination evaluation benchmark with 9,000 questions and 1,500 images from MSCOCO Lin et al.
+(2014), A-OKVQA Schwenk et al.
+(2022), and GQA Hudson and Manning (2019), covering three negative sampling settings (Random, Popular, and Adversarial).
+We also evaluate our method on the CHAIR Rohrbach et al.
+(2018) and AMBER Wang et al.
+(2023a) benchmarks.
+Due to space constraints, we leave the results in Appendix D.
+LVLMs.
+We use 6 representative models to study the effectiveness of CICD, which are InstructBLIP Dai et al.
+(2023), LLaVA-1.5 Liu et al.
+(2024b), LLaVA-Next Liu et al.
+(2024c), Qwen-VL-Chat Bai et al.
+(2023b), Qwen2-VL-Instruct Wang et al.
+(2024a), and Qwen2.5-VL-Instruct Bai et al.
+(2025) respectively.
+The parameter counts of the above models range from 7B to 8B unless otherwise stated.
+Baselines.
+We compare CICD with the following representative training-free methods: VCD Leng et al.
+(2024): a method that introduces noise into the input image during contrastive decoding to suppress language priors.
+ICD Wang et al.
+(2024b): a method that also uses contrastive decoding but adopts adversarial instruction to construct negative logits.
+IBD Zhu et al.
+(2025a): a method that increases the influence of image information during decoding through strengthening the model’s attention to visual features.
+PAI Liu et al.
+(2024d): a method that removes the visual input from the original input and retains the pure textual context as the negative context, while simultaneously enhancing the model’s attention to visual tokens in the original context.
+DeGF Zhang et al.
+(2025): a method that converts the generated caption into an image and re-generates the caption from that image, leveraging the preserved visual content to reduce hallucinations.
+The implementation details of baselines are provided in the Appendix B.
+Implementation Details.
+We use sampling decoding for next-token prediction with default settings.
+Following Leng et al.
+(2024), we incorporate adaptive plausibility constraints into the decoding process.
+For image retrieval, we use the MSCOCO 2014 validation set as $\mathcal{D}_{\textrm{img}}$, which contains 40,504 images.
+All experiments are conducted on a single NVIDIA RTX 3090 GPU.
+Unless otherwise noted, we set the threshold $\gamma=-4$ in all our experiments.
+4.2 Main Results
+Method LLaVA-1.5 InstructBLIP Qwen-VL-Chat
+[…]
+Regular 55.7 17.4 52.60 58.6 18.0 52.99 9.9 7.3 30.73
+VCD Leng et al. (2024) 55.7 16.8 52.91 59.6 18.9 53.20 7.3 5.9 31.28
+ICD Wang et al. (2024b) 53.9 16.6 52.82 55.7 16.7 53.24 5.6 7.0 29.70
+PAI Liu et al. (2024d) 39.4 11.8 53.49 62.6 18.3 53.27 4.6 3.3 29.54
+IBD Zhu et al. (2025a) 54.1 15.8 52.48 56.4 15.5 54.14 4.6 2.9 29.31
+DeGF Zhang et al. (2025) 55.7 16.6 52.72 59.1 17.4 53.06 5.6 3.9 30.21
+CICD (ours) 45.6 13.1 55.80 45.7 13.2 54.20 5.9 3.5 31.34
+Table 2: Results on the COCO subset of DetailCaps. Cs, Ci, and CAP denote CHAIRs, CHAIRi, and CAPTURE respectively.
+Results on POPE.
+Table 1 presents the results of all baseline methods alongside our proposed method.
+Our approach achieves the highest F1 scores across all sampling settings and backbone LVLMs.
+These results suggest that CICD effectively mitigates hallucinations and promotes the use of faithful visual information in LVLMs.
+Notably, under the Popular setting, which uses frequent objects as negative samples and thus increases the likelihood of over-reliance on language priors.
+Nevertheless, our CICD method still achieves significant improvements.
+This highlights its effectiveness in reducing dependency on language priors.
+Results on DetailCaps.
+Table 2 presents the results on the MSCOCO subset of DetailCaps.
+We primarily use CAPTURE as the main evaluation metric, which assesses caption quality through both lexical matching and semantic alignment by measuring the consistency of key content (entities, attributes, and relations) between generated captions and multiple ground-truth references.
+Our CICD method achieves the best performance across all three LVLMs, demonstrating strong image captioning capabilities and confirming its generalizability across different datasets.
+These results suggest that CICD is effective in mitigating the over-reliance on language priors.
+4.3 Comparison of Retrieval Methods
+[…]
+[…]
+Regular 54.6 16.4 72.6 82.25 78.99 77.13
+[…]
+[…]
+Table 3: Comparison of CICD with different contrastive input construction methods on the CHAIR and POPE benchmarks. Ran, Pop, and Adv denote the Random, Popular, and Adversarial subsets, respectively.
+[…]
+We further analyze the similarity between the original and contrastive visual inputs by computing the overlap ratio of image-related words in their descriptions within the COCO dataset.
+We find that the overlap ratios for random sampling and image retrieval are only 0.24% and 0.11%, respectively, relative to the original inputs, which are significantly lower than those of other methods, as shown in Figure 7.
+These findings suggest that using a visually distinct image is effective for contrastive decoding.
+Moreover, random sampling can achieve performance comparable to image retrieval, provided that the dataset $\mathcal{D}_{\textrm{img}}$ contains sufficiently diverse images.
+4.4 Analysis
+Method CHAIR64 CHAIR512
+$downarrow Cs\downarrow Ci\downarrow$
+PAI 20.0 6.2 41.2 10.4
+PAI (w/ DI) 18.0 5.0 40.2 11.2
+VCD 25.0 8.3 59.8 17.8
+VCD (w/ DI) 22.6 7.4 45.0 12.1
+Table 4: CHAIR64 and CHAIR512 represent results on the CHAIR benchmark with maximum generation lengths of 64 and 512, respectively.
+(w/ DI) indicates the method with our proposed distinct contrastive inputs.
+Distorted Contrastive Distributions.
+We analyze the effect of distorted distributions resulting from contrastive visual inputs.
+As previously discussed, methods such as VCD Leng et al.
+(2024) and PAI Liu et al.
+(2024d) introduce a training–inference gap due to the use of noisy or absent visual inputs.
+To examine whether these distorted distributions negatively impact the performance of contrastive decoding, we directly apply our contrastive input construction method to both VCD and PAI.
+The results, shown in Table 4, indicate that the performance of both methods improves when using our construction approach.
+This confirms that distorted contrastive distributions can indeed degrade the effectiveness of contrastive decoding.
+[…]
+[…]
+LLaVA-Next Regular 41.8 11.9 60.2 85.47 83.14 79.87 38.0 10.4 61.06
+CICD 37.6 10.4 62.7 88.22 86.29 82.76 36.2 8.5 62.19
+Qwen2-VL Regular 53.8 10.0 69.4 88.33 85.82 82.43 51.3 10.6 60.04
+CICD 50.6 9.4 73.6 90.67 88.07 84.53 48.4 9.9 61.79
+Qwen2.5-VL Regular 41.2 11.5 65.7 87.83 85.66 83.12 42.2 11.7 62.62
+CICD 39.4 9.6 68.1 89.95 88.00 84.75 43.8 10.9 65.07
+Table 5: Results on additional LVLMs.
+CHAIR512 denotes results on the CHAIR benchmark with a maximum generation length of 512.
+Ran., Pop., and Adv.
+denote the Random, Popular, and Adversarial subsets, respectively.
+Figure 7: Overlap ratio of image-related words between original and negative contexts in contrastive decoding, based on 100 randomly selected COCO images.
+Figure 8: Correlation between visual information overlap ratio and hallucination behavior.
+Visual Information Overlap.
+We analyze the visual information overlap between the original and contrastive contexts across different contrastive decoding methods by computing the overlap ratio of image-related words in their descriptions using the MSCOCO dataset.
+Implementation details can be found in Appendix C.
+Figure 7 presents the overlap ratios for various methods.
+The results show that VCD, ICD, and DeGF exhibit substantial overlap with the original input, indicating that their contrastive visual inputs retain a significant amount of shared visual content.
+In contrast, our method yields a nearly negligible overlap, suggesting it provides more distinct contrastive contexts.
+The information overlap between the two contexts in contrastive decoding confuses the model, amplifying the advantages of suboptimal tokens and thereby inducing hallucinatory behavior.
+Furthermore, we analyze the association between overlap ratios and hallucination behaviors based on the DeGF, which has balanced hallucination rates and overlap ratios.
+We remove samples without hallucinations and without visual information overlap.
+Although the statistical results in Figure 8 do not show a strong correlation ($p=0.076$), there is still a positive association ($\rho=0.29$).
+In short, CICD achieves significantly lower hallucination rates compared to VCD, ICD, and DeGF, demonstrating the superiority of low visual overlap in CICD.
+Method CHAIR64 CHAIR512
+$downarrow Cs\downarrow Ci\downarrow$
+Selective Removal of Language Priors
+CICD ($\gamma=-4$) 18.0 6.1 43.8 11.7
+CICD (w/ DeGF) 21.8 6.9 56.8 16.4
+Complete Removal of Language Priors
+CICD ($\gamma=$-$\infty$) 19.6 6.1 44.4 11.7
+CICD (w/ VCD) 22.6 7.4 45.0 12.1
+Table 6: Comparison among different contrastive decoding methods. CICD (w/ VCD) is equivalent to VCD (w/ DI).
+Excessive Suppression of Language Priors.
+While contrastive decoding is effective in suppressing language priors, it also raises concerns about excessive suppression, which may negatively affect the coherence of generated text.
+To investigate this issue, we set $\gamma$ to -$\infty$ to mitigate all priors and further replace Eq.
+4 with other contrastive decoding methods to evaluate the superiority of dynamic language prior suppression.
+As shown in Table 6, we observe that using cross-image consistency to control the extent of language prior suppression is highly effective, leading to reduced hallucination.
+Moreover, preserving essential priors is crucial, as the performance of CICD degrades when $\gamma$ is set to -$\infty$.
+VCD removes all language priors while DeGF augments the retained language priors, both of which yield suboptimal performance.
+In particular, DeGF’s augmentation of retained priors can harm model performance when prior issues are more severe, such as with longer contexts.
+These results indicate that the dynamic language prior suppression method can effectively mitigate language priors without harming the performance of LVLMs.
+The impact of the choice of $\gamma$ will be discussed in Appendix D.6.
+4.5 Effectiveness of CICD on More LVLMs
+To comprehensively demonstrate the effectiveness and generalizability of our CICD method, we conduct additional experiments using three of the most recent open-source LVLMs.
+The results, presented in Table 5, show that our method consistently achieves significant performance improvements across all experiments.
+CICD effectively reduces hallucinations in both generative and discriminative tasks by mitigating the influence of language priors.
+These results further confirm the transferability of CICD to a variety of LVLMs.
+5 Conclusion
+In this paper, we first reveal that language priors are intrinsic to LVLMs and remain consistent across different images.
+Building on this observation, we propose Cross-Image Contrastive Decoding (CICD), a training-free method that effectively suppresses language priors without compromising the model’s performance.
+Despite its simplicity and efficiency, CICD demonstrates strong practical potential and can be seamlessly integrated with other hallucination mitigation methods to further enhance LVLM performance.
+Acknowledgments
+This work is supported by Zhongguancun Academy Project No.20240103.
+Appendix A Limitation
+We observe that language priors remain consistent across different images.
+Based on this insight, we propose the Cross-Images Contrastive Decoding (CICD) method to eliminate language priors.
+Our method has two main limitations compared to regular decoding.
+First, it requires an additional image that differs sufficiently from the original input.
+Second, both the original and the additional images must undergo separate forward passes to obtain their respective logit distributions for contrastive decoding.
+Addtional Image
+Although the additional image is indispensable, its acquisition is straightforward.
+Hallucinations typically occur in content-rich elements such as objects, attributes, and relationships, which naturally exhibit substantial variability across real-world images.
+We demonstrate that random selection and retrieval-based methods achieve comparable low visual information overlap ratios when selecting images.
+The experimental results also show that randomly selected images perform comparably to retrieved ones, highlighting the practical applicability of our method.
+This flexibility significantly lowers the barrier for real-world deployment.
+Twice Forward Propagation
+Although performing twice forward propagations doubles the theoretical computational cost, in practice, we merge the two inputs into a single batch and perform one forward propagation with a doubled batch size to achieve the same effect.
+Under this implementation, our method incurs only a 10% increase in inference time and a 5% increase in memory usage, demonstrating strong potential for practical deployment.
 MaxLen Method LLaVA-1.5 InstructBLIP
@@
 64 Regular 24.4 8.9 56.6 53.7 35.6 13.2 56.4 54.8
-VCD 25.0 8.3 59 54.0 32.2 10.3 60.6 55.2
+VCD 25.0 8.3 59.0 54.0 32.2 10.3 60.6 55.2
 ICD 23.2 8.1 58.4 54.4 29.8 9.8 60.6 55.7
 IBD 21.2 6.9 58.8 54.5 27.8 9.2 60.3 56.1
-DeFG 22.4 7.2 58.2 54.4 32.4 11 59.4 56.0
+PAI 20.0 6.2 56.9 55.1 26.0 8.9 53.9 54.4
+DeGF 22.4 7.2 58.2 54.4 32.4 11 59.4 56.0
 CICD 18.0 6.1 59.6 53.3 23.8 7.7 62.2 54.8
@@
 IBD 57.6 16.5 74.2 104.3 57.6 15.7 70.8 118.5
-DeFG 57.4 16.3 75.5 102.8 59.0 17.7 71.5 108.9
+PAI 41.2 10.4 68.6 116.2 67.6 19.4 68.0 143.1
+DeGF 57.4 16.3 75.5 102.8 59.0 17.7 71.5 108.9
 CICD 43.8 11.7 75.0 103.3 49.8 13.7 70.3 98.8
@@
 IBD 9.8 50.5 42.2 4.4 69.2 86.3 62.1 72.2 81.20
-DeFG 9.1 50.7 39.9 4.1 70.2 86.8 63.0 73.0 81.95
+PAI 7.7 49.3 36.9 3.3 70.5 86.5 64.2 73.7 83.0
+DeGF 9.1 50.7 39.9 4.1 70.2 86.8 63.0 73.0 81.95
 CICD 6.6 52.7 34.8 2.2 71.1 89.4 61.9 73.1 83.25
@@
 IBD 9.0 56.1 45.1 4.6 54.2 86.3 34.2 49.0 70.00
-DeFG 9.7 54.1 44.5 5.2 72.1 80.7 75.1 77.8 84.05
+PAI 11.7 52.8 55.1 5.4 68.0 78.3 67.0 72.2 80.25
+DeGF 9.7 54.1 44.5 5.2 72.1 80.7 75.1 77.8 84.05
 CICD 7.1 53.6 35.0 2.3 72.8 85.4 70.1 77.0 84.95
-Table 3: Results on AMBER. The maximum generation length is 512. AMBER metirc is calculated by $(100-CHAIR+F1)/2$.
-5.1 Experiment Settings
-Due to space constraints, we present only the key aspects of our experimental setup here. Detailed settings can be found in Sec.B.
-Benchmarks
-We evaluate the performance of CICD across 4 widely adopted multimodal hallucination benchmarks to validate its effectiveness on different tasks.
-These include two generative benchmarks (CHAIR Rohrbach et al.
-(2018) and DetailCaps Ye et al.
-(2025)), one discriminative benchmark (POPE Li et al.
-(2023b)), and one hybrid benchmark (AMBER Wang et al.
-(2024a)).
-Evaluated LVLMs
-We evaluate the generalizability of CICD across 6 LVLMs from three model families: InstructBLIP Dai et al.
-(2023), the LLaVA series (LLaVA-1.5 Liu et al.
-(2024c) and LLaVA-Next Liu et al.
-(2024d)), and the Qwen series (Qwen-VL-Chat Bai et al.
-(2023b), Qwen2-VL-Instruct Wang et al.
-(2024b), and Qwen2.5-VL-Instruct Bai et al.
-(2025)).
-Unless stated otherwise, all models are implemented at the 7B scale.
-Baselines
-We compare CICD with 4 recent training-free methods that specifically target language priors: VCD Leng et al.
-(2024), ICD Wang et al.
-(2024c), IBD Zhu et al.
-(2024b), and DeFG Zhang et al.
-(2025).
-All of these methods construct negative contexts based on the original image.
-In contrast, we identify several limitations of this approach and propose CICD as a more effective alternative.
-5.2 Implementation Details
-We use sampling decoding for next-token prediction with default settings.
-Following Leng et al.
-(2024), we incorporate adaptive plausibility constraints into the decoding process.
-We use the MSCOCO 2014 validation set as the image retrieval corpus, which contains 40,504 images.
-All experiments are conducted on an NVIDIA RTX 3090 GPU.
-5.3 Main Results
-Method LLaVA-1.5 InstructBLIP Qwen-VL-Chat
+Table 8: Results on AMBER. The maximum generation length is 512. AMBER metirc is calculated by (100CHAIR+F1)/2(100-CHAIR+F1)/2.
+Appendix B Implementation Details for Baselines
+The implementation details for the reproduction of baselines are as follows:
+  • VCD Leng et al.
+(2024): We added 500 steps of Gaussian noise to the images to construct the negative visual input for contrastive decoding.
+The hyperparameters in VCD are set to $\alpha=1,\beta=0.1$.
+  • ICD Wang et al.
+(2024b): We use the sentence "You are a confused objects detector to provide a fuzzy overview or impression of the image." as the disturbance instruction for contrastive decoding.
+The hyperparameters for contrastive decoding are set according to VCD.
+  • IBD Zhu et al. (2025a): We follow the practice in the original paper, setting $\epsilon=2$, $\alpha=1.5\times 10^{4}$, and $\beta=0.1$ in our experiments.
+  • PAI Liu et al.
+(2024d): We set the visual attention augmentation coefficient $\alpha$ to 0.5 and enable increased visual attention starting from layer 2 (counting from 0).
+The hyperparameter $\gamma$ is set to 1.1.
+PAI employs adaptive plausibility constraints with $\beta=0.1$.
+  • DeGF Zhang et al.
+(2025) We employ LLaVA-v1.5-7B to generate captions for each image and use Stable Diffusion v1-5 to generate images from the captions.
+The parameter $\gamma$ is set to 0.1, and $\beta$ is set to 0.1 for the image captioning task and 0.25 for the visual question answering task.
+Note: The same parameters may have different meanings across methods. For specific details, please refer to the original papers.
+Appendix C Implementation Details for Visual Information Overlap
+First, we apply several contrastive decoding methods, each constructing its own negative context, and use them to prompt LVLMs to answer user queries.
+For comparison, we obtain standard responses via regular decoding.
+We then use QwenVL-Max-0125 to extract image-related words from the generated responses.
+Inspired by CoT Wei et al.
+(2022), we guide the model through a multi-stage extraction process: it first identifies all visually relevant words without access to the image, and then selects those strongly associated with the image content, which we define as relevant words.
+Finally, we compute the overlap ratio between the relevant words generated from negative contexts and those obtained through regular decoding.
+This ratio quantifies how much essential visual content is preserved in responses generated from negative contexts, serving as an indicator of retained visual information.
+The prompt used to guide QwenVL-Max is shown in Figure 10.
 […]
-Regular 55.7 17.4 52.60 58.6 18.0 52.99 9.9 7.3 30.73
-VCD 55.7 16.8 52.91 59.6 18.9 53.20 7.3 5.9 31.28
-ICD 53.9 16.6 52.82 55.7 16.7 53.24 5.6 7.0 29.70
-IBD 54.1 15.8 52.48 56.4 15.5 54.14 4.6 2.9 29.31
-DeFG 55.7 16.6 52.72 59.1 17.4 53.06 5.6 3.9 30.21
-CICD 45.6 13.1 55.80 45.7 13.2 54.20 5.9 3.5 31.34
-Table 4: Results on the COCO subset of DetailCaps. The maximum generation length is 512. Cs, Ci, and CAP represent CHAIRs, CHAIRi, and CAPTURE respectively.
-Results on POPE
-As shown in Tab.
-1, our method achieves the highest F1 scores across all experimental settings.
-This suggests that we can utilize cross-image consistency to effectively alleviate language priors and promote LVLMs to convey faithful visual information.
-The Popular setting uses frequent objects for negative samples, increasing susceptibility to language priors.
-The strong performance gains of our method in this setting highlight its effectiveness in mitigating such priors.
-Model Method CHAIR(MaxLen=512) POPE (F1$\uparrow$)
 […]
-LLaVA-Next Regular 41.8 11.9 60.2 85.47 83.14 79.87
-CICD 37.6 10.4 62.7 88.22 86.29 82.76
-Qwen2-VL-Instruct Regular 53.8 10.0 69.4 88.33 85.82 82.43
-CICD 50.6 9.4 73.6 90.67 88.07 84.53
-Qwen2.5-VL-Instruct Regular 41.2 11.5 65.7 87.83 85.66 83.12
-CICD 39.4 9.6 68.1 89.95 88.00 84.75
-Table 5: Results on more LVLMs. F1 is the average F1-score across three datasets (MSCOCO, A-OKVQA, and GQA).
-Results on CHAIR
-The results in Tab.
-2 demonstrate our method’s superior performance across all experimental conditions, with particularly pronounced improvements observed at longer generation lengths.
-Longer context lengths amplify the model’s susceptibility to language priors, making these experimental results particularly compelling evidence of our method’s efficacy in mitigating such priors.
-Compared to baselines, our method’s principal advantage lies in its complete preservation of visual information and essential language priors, which is directly reflected in the superior performance in image captioning.
-Results on AMBER
-As illustrated in Tab.
-3, our method achieves the lowest hallucination frequency in generative tasks while simultaneously attaining state-of-the-art performance on the comprehensive AMBER metric.
-Language prior is particularly critical for generative tasks, where language priors manifest more severely than in discriminative tasks, explaining our method’s pronounced advantages in image captioning.
-Results on DetailCaps
-Tab.
-4 presents the evaluation results on DetailCaps.
-The CAPTURE metric quantitatively evaluates caption quality through both lexical matching and semantic alignment, measuring the consistency of key content (entities, attributes, and relations) between generated captions and multiple ground-truth references.
-Our method achieves the best performance across all three evaluated LVLMs, demonstrating its robust capability to mitigate language priors and significantly enhance image captioning performance.
-5.4 Effectiveness on More LVLMs
-To comprehensively demonstrate the effectiveness and generalizability of our method, we select three of the most recent open-source LVLMs to conduct experiments.
-The results are presented in Tab.
-5.
-Our method demonstrates significant performance improvements in all experiments, consistently reducing hallucinations across both generative and discriminative tasks through effective mitigation of language priors.
-We also test our method on these LVLMs by DetailCaps and the results are shown in Tab.
-6.
-The improvements observed across different LVLMs confirm our method’s practical advantages: an elegantly simple framework coupled with state-of-the-art effectiveness.
-Further experiments will be discussed in Sec. B.
-6 Conclusion
-In this paper, we mitigate the influence of language priors in LVLMs.
-Through a comprehensive investigation, we reveal that language priors are intrinsic to the model and remain consistent across different images.
-Building on this observation, we propose Cross-Image Contrastive Decoding (CICD) — a training-free method that effectively mitigates language priors while preserving visual information.
-Extensive experiments across multiple benchmarks and LVLM architectures demonstrate the effectiveness of CICD, particularly in image captioning tasks where language priors are most prominent.
-The simplicity and efficiency of CICD offer strong practical potential, enabling seamless integration with other hallucination mitigation methods to further enhance LVLM performance.
-This integration forms a key direction for our future work.
-7 Limitation
-We observe that language priors remain consistent across different images.
-Based on this insight, we propose the Cross-Images Contrastive Decoding (CICD) method to eliminate language priors.
-Our method has two main limitations compared to regular decoding.
-First, it requires an additional image that differs sufficiently from the original input.
-Second, both the original and the additional images must undergo separate forward passes to obtain their respective logit distributions for contrastive decoding.
-Addtional Image
-While the additional image is necessary, there is no strict requirement on its degree of difference from the original.
-Hallucinations typically occur in content-rich elements such as objects, attributes, and relationships, which naturally exhibit substantial variability across real-world images.
-Furthermore, our experiments demonstrate that the effectiveness of hallucination mitigation is not strongly correlated with image similarity.
-The experimental results demonstrate that randomly selected images perform comparably to retrieved ones, highlighting the practical applicability of our method.
-This flexibility significantly lowers the barrier for real-world deployment.
-Twice Forward Propagation
-Although performing twice forward propagations doubles the theoretical computational cost, in practice, we merge the two inputs into a single batch and perform one forward propagation with a doubled batch size to achieve the same effect.
-Under this implementation, our method incurs only a 10% increase in inference time and a 5% increase in memory usage, demonstrating strong potential for practical deployment.
-Appendix A Detailed Related Work
-A.1 Large Vision-Language Models
-Typically, Large Vision-Language Models (LVLMs) Dai et al.
-(2023); Liu et al.
-(2024c); Bai et al.
-(2023b); Zhu et al.
-(2024a) consist of a visual encoder Radford et al.
-(2021); Dosovitskiy et al.
-(2021) and a Large Language Model (LLM) backbone Touvron et al.
-(2023); Chiang et al.
-(2023).
-The visual encoder extracts visual features, which are then aligned with text tokens through a fusion module to form the input context for the LLM.
-Through multi-stage training, including modality alignment and instruction tuning Liu et al.
-(2023), LVLMs acquire the ability to perform various multimodal tasks such as image captioning Ye et al.
-(2025), visual question answering Yu et al.
-(2024), and visual grounding Lu et al.
-(2024).
-Despite significant advancements, LVLMs still suffer from severe hallucination problems, often generating content that contradicts the visual input.
-A.2 Hallucination in LVLMs
-Originally, hallucinations Ji et al.
-(2023); Liu et al.
-(2024b) were defined as undesirable outputs or non-factual information.
-In the multimodal domain, the study of hallucinations primarily focuses on cross-modal consistency Li et al.
-(2023c); Wang et al.
-(2023), requiring that LVLM-generated content remains faithful to the visual inputs.
-Multiple factors contribute to hallucinations in LVLMs, including biased visual attention Zhu et al.
-(2025); Li et al.
-(2025), attention misallocation Huang et al.
-(2024); Yin et al.
-(2025); Kang et al.
-(2025), and language priors Leng et al.
-(2024); Zhu et al.
-(2024b); Wang et al.
-(2024c).
-Among these, language priors are the most direct cause, making it difficult for LVLMs to accurately express their understanding of visual information.
-A.3 Language Priors
-As LVLMs are built upon LLMs, they inherently inherit the linguistic knowledge embedded in LLMs, which are referred to as language priors Zhibo et al.
-(2023); Wu et al.
-(2022).
-Language priors specifically manifest when LVLMs generate content that is linguistically plausible but inconsistent with visual evidence.
-Some works Chen et al.
-(2025); Liu et al.
-(2024a); Gunjal et al.
-(2024) meticulously design specialized training data to mitigate this issue.
-However, such training-based approaches exhibit limited generalizability and cannot be easily transferred to broader application scenarios.
-Recently, a series of training-free approaches have directly reduced language priors using contrastive decoding Li et al.
-(2023a); Chuang et al.
-(2024), achieving remarkable performance.
-These methods construct an alternative logit distribution on top of the original one through techniques such as masking the image Leng et al.
-(2024), perturbing the instruction Wang et al.
-(2024c), augmenting the vision input Zhu et al.
-(2024b), or performing cross-modal conversion Zhang et al.
-(2025).
-During decoding, the two logit distributions are contrasted to eliminate language priors.
-While these methods effectively mitigate linguistic priors, they inevitably incur visual information loss during contrastive decoding due to information overlap, where both distributions share partial visual features.
-Through a comprehensive analysis of language priors, we identify their vision-independent nature.
-Based on this insight, we propose the Cross-Images Contrastive Decoding (CICD) method, which eliminates language priors by leveraging their consistency across different images, while preserving visual information through inter-image variation.
-Appendix B Detailed Experiment Settings
-B.1 Benchmarks
+LLaVA-1.5-13B Regular 58.8 17.0 73.4 82.94 81.46 78.49
+CICD 44.4 12.9 76.0 87.92 85.68 80.99
+INstructBLIP-13B Regular 65.6 21.0 62.7 81.31 77.20 74.69
+CICD 44.0 11.6 65.7 87.52 83.60 78.96
+Table 9: Results on larger LVLMs. The maximum generation length is 512.
+Appendix D More Experimental Results
+D.1 Benchmarks
 In order to validate the effectiveness of our method in mitigating language hallucination, we evaluate its performance on four widely-used multimodal hallucination benchmarks: two generative benchmarks (CHAIR Rohrbach et al.
@@
 (2023a)).
-Method LLaVA-Next Qwen2-VL-Instruct Qwen2.5-VL-Instruct
-[…]
-Regular 61.06 38.0 10.4 60.04 51.3 10.6 62.62 42.2 11.7
-CICD 62.19 36.2 8.5 61.79 48.4 9.9 65.07 43.8 10.9
-Table 6: Results on the COCO subset of DetailCaps with more LVLMs.
 CHAIR
@@
 CHAIRs $=\frac{|\text{Hallucinated Objects}|}{|\text{All Objects}|},$ (6)
-CHAIRi $=\frac{|\text{Hallucinated Sentences}|}{|\text{All Sentences}|}$
+CHAIRi $=\frac{|\text{Hallucinated Captions}|}{|\text{All Captions}|}$
 DetailCaps
-is a fine-grained image captioning benchmark, accompanied by ground-truth detail captions generated by GPT-4V Yang et al.
+is a fine-grained image captioning benchmark, accompanied by ground-truth detail captions generated by GPT-4V Wu et al.
 (2024), Gemini-1.5-Pro Team et al.
-(2024), and GPT-4o OpenAI et al.
+(2024), and GPT-4o Hurst et al.
 (2024) for evaluation.
@@
 Cover $=\frac{\mathrm{len}(R^{\prime}_{obj}\cap A_{obj})}{\mathrm{len}(A_{obj})},$
-Hal =$=\frac{\{\text{CHAIR}>0\}}{\{\text{All Caps}\}},$
+Hal $=\frac{\{\text{CHAIR}>0\}}{\{\text{All Caps}\}},$
 Cog $=\frac{\mathrm{len}(R^{\prime}_{obj}\cap H_{obj})}{\mathrm{len}(R^{\prime}_{obj})},$
 where $H_{obj}$ denotes the set of hallucinated target objects generated by the LVLMs, and All Caps refers to all generated captions.
-B.2 Evaluated LVLMs
+D.2 Evaluated LVLMs
 To demonstrate the generalizability of the proposed CICD as a broadly applicable plug-and-play module, we evaluate its performance across six different LVLMs spanning three model families: InstructBLIP Dai et al.
 (2023), LLaVA-1.5 Liu et al.
-(2024c), LLaVA-Next Liu et al.
-(2024d), Qwen-VL-Chat Bai et al.
+(2024b), LLaVA-Next Liu et al.
+(2024c), Qwen-VL-Chat Bai et al.
 (2023b), Qwen2-VL-Instruct Wang et al.
-(2024b), and Qwen2.5-VL-Instruct Bai et al.
+(2024a), and Qwen2.5-VL-Instruct Bai et al.
 (2025).
 InstructBLIP and LLaVA-1.5 are both built upon Vicuna-7B Chiang et al.
-(2023) as their language backbone, while LLaVA-Next is based on LLaMA3.
+(2023b) as their language backbone, while LLaVA-Next is based on LLaMA3.
 The Qwen-VL series, including Qwen2-VL-Instruct and Qwen2.5-VL-Instruct, is developed using the QwenLM framework.
@@
 All models are implemented at the 8B parameter scale.
-Model Method CHAIR(MaxLen=512) POPE (F1$\uparrow$)
-[…]
-LLaVA-1.5-13B Regular 58.8 17.0 73.4 82.94 81.46 78.49
-CICD 44.4 12.9 76.0 87.92 85.68 80.99
-INstructBLIP-13B Regular 65.6 21.0 62.7 81.31 77.20 74.69
-CICD 44.0 11.6 65.7 87.52 83.60 78.96
-Table 7: Results on larger LVLMs. The maximum generation length is 512.
-B.3 Baselines
-We compare our method with several training-free baselines that specifically target language priors.
-VCD Leng et al. (2024) introduces noise into the input image to suppress language priors, and then applies contrastive decoding for their elimination.
-ICD Wang et al. (2024c) constructs negative logits using adversarial instructions and enhances positive signals through contrastive decoding.
-IBD Zhu et al. (2024b) strengthens the model’s attention to visual features, thereby increasing the influence of image information during decoding.
-DeFG Zhang et al. (2025) converts the generated caption into an image and re-generates the caption from that image, leveraging the preserved visual content to reduce hallucinations.
-Appendix C More Experiment Results
-Figure 7: The correlation between similarity and performance.
-We conducted this experiment on the validation set of CHAIR.
-Sim-CLIP represents normalized image similarity, measured as the cosine distance between the features of two images extracted by CLIP.
-C.1 Effectiveness on Larger LVLMs
-We evaluate the effectiveness of our method on larger models using two benchmarks, POPE and CHAIR, with the results presented in Tab.7.
+D.3 Results on CHAIR
+The results in Table 7 demonstrate our method’s superior performance across all experimental conditions, with particularly pronounced improvements observed at a longer generation length.
+Longer context lengths amplify the model’s susceptibility to language priors, making these experimental results particularly compelling evidence of our method’s efficacy in mitigating such priors.
+Compared to baselines, our method’s principal advantage lies in its complete preservation of visual information and essential language priors, which is directly reflected in the superior performance in image captioning.
+D.4 Results on AMBER
+As illustrated in Table 8, our method achieves the lowest hallucination frequency in generative tasks while simultaneously attaining state-of-the-art performance on the comprehensive AMBER metric.
+Language prior is particularly critical for generative tasks, where language priors manifest more severely than in discriminative tasks, explaining our method’s pronounced advantages in image captioning.
+D.5 Effectiveness on Larger LVLMs
+We evaluate the effectiveness of our method on larger models using two benchmarks, POPE and CHAIR, with the results presented in Table9.
 Our method continues to deliver significant performance improvements on a 13B model, indicating its strong generalizability.
 These results suggest that our approach scales well and remains effective across different model sizes.
-C.2 Divergence Between Images
-Contrastive decoding is a method that suppresses consistent features between two distributions while amplifying distinctive ones.
-To prevent the loss of visual information during contrastive decoding, we select an image with sufficient discrepancy to construct a contrasting distribution.
-Thus, while maintaining language prior consistency, the divergent visual information helps avoid erroneous debiasing caused by feature overlap.
-We use CLIP to retrieve images based on cosine similarity, and select the image with the lowest similarity as the contrastive image.
-We conduct experiments to explore the correlation between similarity and performance, as shown in Fig.
-7.
-We observe that the similarity between images exhibits no strong correlation with performance.
-Compared to regular decoding methods, images with varying similarity levels all demonstrate significant improvements.
-As discussed in Sec.4.1, the influence of language priors is primarily reflected in the LVLMs’ comprehension of objects, attributes, and relationships.
-This information exhibits limited overlap across real-world images.
-We hypothesize that random image selection could achieve comparable effectiveness, with experimental results presented in Tab.8.
-Experimental results demonstrate that random image selection can also significantly mitigate language priors in LVLMs.
-In practical applications, the random selection approach offers greater convenience, substantially enhancing the usability of our method.
-In Tab.
-8, we also attempted to construct language priors by removing visual inputs, but found this approach less effective compared to eliminating linguistic priors through cross-image consistency.
-We attribute this limitation to LVLMs being trained on large-scale multimodal data, where removing visual inputs disrupts the model’s intrinsic representations and leads to information distortion.
-Consequently, language priors in LVLMs cannot be directly constructed, which further substantiates the necessity of leveraging cross-image consistency in our method.
-[…]
-CHAIR(MaxLen=512)
-Regular 54.6 16.4 72.6
-w/o Image 54.6 15.5 76.6
-Random Image 46.6 13.2 76.7
-Retrieved Image 43.8 11.7 75.0
-CHAIR(MaxLen=64)
-Regular 24.4 8.9 56.6
-w/o Image 25.4 7.6 61.9
-Random Image 21.6 6.9 58.8
-Retrieved Image 18.0 6.1 59.6
-Table 8: This experiment was conducted by LLaVA-1.5.
-Random Image and Retrieved Image represent the contrastive images selected through random sampling and image retrieval, respectively.
-w/o Image represents the contrastive distribution that is constructed by removing the input image.
-C.3 Threshold of Detrimental Priors
+D.6 Effects of $\gamma$
 While language priors contribute to hallucinations in LVLMs, they also play a critical role in maintaining textual coherence and fluency.
-We employ JS divergence to distinguish between essential and detrimental priors, where essential priors exhibit high cross-image consistency, reflected in lower JS divergence scores.
-We establish a threshold of $\gamma=-4$ in Eq.6, based on the statistical distribution of JS divergence scores shown in Fig.6.
-A comprehensive analysis is shown in Fig.
-8.
+We employ JS divergence to selectively suppress language priors, where priors exhibiting high cross-image consistency are retained, as reflected by lower JS divergence scores.
+We establish a threshold of $\gamma=-4$ in Eq.
+2, based on the statistical distribution of JS divergence scores.
+A comprehensive analysis is shown in Figure 9.
 The experimental results align with the statistical analysis, demonstrating optimal performance at $\gamma=-4$.
-Another notable finding is that preserving essential priors proves meaningful, as it ensures fluent and coherent generation contexts.
-Appendix D Analysis of Visual Information Overlap
-Fig. 2 presents the overlap ratio of image-related words between responses generated from negative contexts and those from regular decoding. Below, we detail the analysis procedure.
-First, we apply several contrastive decoding methods, each constructing its own negative context, and use them to prompt LVLMs to answer user queries.
-For comparison, we obtain standard responses via regular decoding.
-We then use QwenVL-Max-0125 Bai et al.
-(2023b) to extract image-related words from the generated responses.
-Inspired by CoT Wei et al.
-(2022), we guide the model through a multi-stage extraction process: it first identifies all visually relevant words without access to the image, and then selects those strongly associated with the image content, which we define as relevant words.
-Finally, we compute the overlap ratio between the relevant words generated from negative contexts and those obtained through regular decoding.
-This ratio quantifies how much essential visual content is preserved in responses generated from negative contexts, serving as an indicator of retained visual information.
-The prompt used to guide QwenVL-Max is shown in Fig.
-9.
-Based on this analysis, Fig.
-2 shows that VCD and ICD retain over 30% of essential visual information, which is substantially weakened during contrastive decoding.
-In contrast, our method results in a negligible amount of residual information, effectively preventing visual information loss.
-Figure 8: The results on the validation set of CHAIR using LLaVA-1.5. The X-axis represents the value of $\gamma$.
+Another notable finding is that preserving necessary priors is beneficial, as it prevents the model’s generative capability from being compromised during the contrastive decoding process, thereby further reducing hallucinations.
+Figure 9: The results on the validation set of CHAIR using LLaVA-1.5. The X-axis represents the value of $\gamma$. $\gamma=$-$\infty$ is equivalent to removing both detrimental and essential priors.
+Figure 10: Prompt for guiding QwenVL-Max to progressively extract relevant words from the response.
+Figure 11: Case study (1).
+Figure 12: Case study (2).
 Appendix E Case Study
-We present case studies to illustrate the superior performance of our method in hallucination reduction, as shown in Fig.
-10 and Fig.
-11.
+We present case studies to illustrate the superior performance of our method in hallucination reduction, as shown in Figure 11 and Figure 12.
 The image captions are generated by LLaVA-1.5.
@@
 We can clearly observe that Regular Decoding produces a substantial amount of hallucinated content, often manifesting as content words such as objects, attributes, and relationships.
-In Fig.
-10, for example, although the image depicts a skateboarding scene, the caption includes a reference to "spectators"—a detail not present in the image.
-Similarly, Fig.
-11 shows an eating scene, yet the caption introduces a "long dining table," which is not actually visible.
+In Figure 11, for example, although the image depicts a skateboarding scene, the caption includes a reference to "spectators"—a detail not present in the image.
+Similarly, Figure 12 shows an eating scene, yet the caption introduces a "long dining table," which is not actually visible.
 These hallucinations stem from LVLMs inheriting prior knowledge from LLMs, leading to biased descriptions in certain typical scenarios.
 In contrast, our method significantly reduces such hallucinations.
-While Fig.
-11 still contains a hallucinated object (chips), this instance can be attributed to image ambiguity, where even human observers might plausibly infer such content.
 Overall, our method effectively mitigates the influence of language priors by leveraging cross-image consistency.
-Figure 9: Prompt for guiding QwenVL-Max to progressively extract relevant words from the response.
-Figure 10: Case study (1).
-Figure 11: Case study (2).